### v1

# The commutativity of cofinal types

Abstract.
We developed the theory of deterministic ideals and present a systematic study of the pseudo-intersection property with respect to an ideal introduced in [3]. We apply this theory to prove that for any two ultrafilters $U,V$ on $\omega$, $U\cdot V\equiv_{T}V\cdot U$. Additionally, we analyze the class of ultrafilters which are Tukey above $\omega^{\omega}$. Our theory applies the study of the Tukey types of a general sum of ultrafilters, which, as evidenced by the results of this paper, can be quite complex. Finally, we introduce the hierarchy of $\alpha$-almost rapid ultrafilters. We prove that it is consistent for them to form a strictly wider class than the rapid ultrafilters, and give an example of a non-rapid $p$-point which is Tukey above $\omega^{\omega}$. This answers several questions from [2, 3].
The research of the first author was supported by the National Science Foundation under Grant No. DMS-2246703
0. Introduction
The Tukey order stands out as one of the most studied orders of ultrafilters [28, 14, 25, 10, 32, 2]. Its origins lie in the examination of Moore-Smith convergence, and it holds particular significance in unraveling the cofinal structure of the partial order $(U,\supseteq)$ of an ultrafilter. Formally, given two posets, $(P,\leq_{P})$ and $(Q,\leq_{Q})$ we say that $(P,\leq_{P})\leq_{T}(Q,\leq_{Q})$ if there is map $f:Q\rightarrow P$, which is cofinal, namely, $f^{\prime\prime}\mathcal{B}$ is cofinal in $P$ whenever $\mathcal{B}\subseteq Q$ is cofinal. Schmidt [33] observed that this is equivalent to having a map $f:P\rightarrow Q$, which is unbounded, namely, $f^{\prime\prime}\mathcal{A}$ is unbounded in $Q$ whenever $\mathcal{A}\subseteq P$ is unbounded in $P$. We say that $P$ and $Q$ are Tukey equivalent, and write $P\equiv_{T}Q$, if $P\leq_{T}Q$ and $Q\leq_{T}P$; the equivalence class $[P]_{T}$ is called the Tukey type or cofinal type of $P$.
The scope of the study of cofinal types of ultrafilters covers several long-standing open problems such as:
  • Isbell’s problem[19]: Is it provable within ZFC that a non-Tukey-top ultrafilter [^1] A Tukey-top ultrafilter is an ultrafilter which is Tukey equivalent to the poset $([\mathfrak{c}]^{<\omega},\subseteq)$. Isbell [19] constructed such ultrafilters in $ZFC$. on $\omega$ exists?
  • Kunen’s Problem: Is it consistent that $\mathfrak{u}_{\aleph_{1}}<2^{\aleph_{1}}$? Namely, is it consistent to have a set $\mathcal{B}\subseteq P(\omega_{1})$ of cardinality less than $2^{\aleph_{1}}$ which generates an ultrafilter?
The Tukey order is also related to the Katovich problem. A systematic study of the Tukey order on ultrafilter on $\omega$ traces back to Isbell [19], later to Milovich [28] and Dobrinen and Todorcevic [14]. Lately, Dobrinen and Benhamou [2] extended this study to ultrafilters on cardinals greater than $\omega$. They found a connection to a parallel recent line of research at the level of measurable cardinals, known as the Galvin property studied by Abraham, Benhamou, Garti, Goldberg, Gitik, Hayut, Magidor, Poveda, Shelah and others [1, 15, 16, 6, 4, 5, 8, 7, 17, 9]. The Galvin property in one of its forms is equivalent to being Tukey top as shown essentially by Isbell (in different terminology). Moreover, being Tukey-top in the restricted class of $\sigma$-complete ultrafilters takes the usual studied forms of the Galvin property.
In this paper, we address the problem of commutativity of the Tukey types of Fubini products of ultrafilters $U,V$ on $\omega$ (Definition 1.1), denoted by $U\cdot V$. This problem was suggested in [3], and was already partially addressed:
  • Dobrinen and Todorchevic [14] proved that if $U,V$ are rapid $p$-points then $U\cdot V\equiv_{T}V\cdot U$.
  • Milovich [29] extended this result to prove that if $U,V$ are just $p$-points, then $U\cdot V\equiv_{T}V\cdot U$.
  • Benhamou and Dobrinen proved later that if $U,V$ are $\kappa$-complete ultrafilters over a measurable cardinal $\kappa$ then $U\cdot V\equiv_{T}V\cdot U$.
The main result of this paper is:
Theorem.
For any ultrafilters $U_{1},U_{2},...,U_{n}$ on $\omega$, and any permutation $\sigma:\{1,...,n\}\rightarrow\{1,...,n\}$, $U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{\sigma(1)}\cdot...\cdot U_{\sigma(n)}$.
To do this we will, in fact, prove that
(*$(*)\ \ \ U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times U_{2}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
from which the commutativity follows straightforwardly.
Our main result stands in sharp contrast to the Rudin-Keisler ordering which is known not to be commutative with respect to Fubini product [^2] For example if $U,W$ are non-isomorphic Ramsey ultrafilters then $U\cdot W\not\equiv_{RK}W\cdot U$. Just otherwise, by a theorem of Rudin (see for example [20, Thm. 5.5], $U,W$ should be Rudin-Frolík (and therefore Rudin-Keisler) comparable, contradicting the RK-minimality of Ramsey ultrafilters.. On measurable cardinals, the situation is even more dramatic, due to a theorem of Solovey (see [20, Thm. 5.7]) if $U,W$ are $\kappa$-complete ultrafilters on $\kappa$ the $U\cdot W\equiv_{T}W\cdot U$ if and only if $W\equiv_{RK}U^{n}$ for some $n$ or vise versa. Recently, Goldberg [18] examined situations of commutativity with respect to several product operations of countable complete ultrafilters.
As a corollary of our main theorem, we conclude the following Tukey-structural results:
Theorem.
The class of ultrafilters $U$ such that $U\cdot U\equiv_{T}U$ is upward closed with respect to the Tukey order.
The main idea that is used in the proof of the main result, as suggested by the equivalence (*)(*)( * ), is to analyze the cofinal types of ideals and filters connected to a given ultrafilter $U$. More specifically, we will exploit the idea of the pseudo-intersection property with respect to $I$ (Definition 2.2) which was introduced in [3] and was used to prove that Milliken-Taylor ultrafilters and generic ultrafilters [^3] For the definition of $I^{\otimes\alpha}$, see the paragraph before Fact 2.11. for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfy $U\cdot U\equiv_{T}U$. In  §2, we provide a comprehensive study of this property. The main result of this section is
Theorem.
Suppose that $\mathcal{A}$ is a discrete set of ultrafilters. Then for each $U\in\mathcal{A}$, $U$ has the $(\cap\mathcal{A})^{*}$-p.i.p.
The property of ideals which will be in frequent use, is the property of being deterministic (Definition 2.22). This property guarantees that whenever $I\subseteq J$, $I\leq_{T}J$.
We also investigate the Tukey type of ultrafilters of the form $\sum_{U}V_{\alpha}$. The only general result regarding the Tukey-type of such ultrafilters is due [^4] Dobrinen and Todorcevic proved in for $\kappa=\omega$, but the proof for a general $\kappa$ appears in [2]. to Dobrinen and Todorcevic [14] where they prove that if $U$ is an ultrafilter on $\kappa$, and $V_{\alpha}$ is a sequence of ultrafilters, then $\sum_{U}V_{\alpha}\leq_{T}U\times\prod_{\alpha<\kappa}V_{\alpha}$. It turns out that the Tukey class of such ultrafilters is much more complicated and the nice characterization we have for $U\cdot V\equiv_{T}U\times V^{\omega}$ is missing in the general case. It is not hard to see that the ultrafilter $\sum_{U}V_{\alpha}$ is Tukey below each ultrafilter in the set $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\kappa{\rangle})=\{U\times\prod_{\alpha\in X}V_{\alpha}\mid X\in U\}$. We prove that in some sense it is the greatest lower bound:
Theorem.
For complete directed ordered set $\mathbb{P}$, $\mathbb{P}$ is uniformly bellow [^5] See definition $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\kappa{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
By putting more assumptions on the sequence of ultrafilters, we are able to get nicer results. One of them is that under the assumption that
$V_{0}\leq_{T}V_{1}\leq_{T}....$
in which case we have that $\sum_{U}V_{n}\equiv_{T}U\times\prod_{n<\omega}V_{n}$. From this special case, we can recover the formula $U\cdot V\equiv_{T}U\times V^{\omega}$. Another set of assumptions we consider is when for each $n<\omega$, $V_{n}\cdot V_{n}\equiv_{T}V_{n}\geq_{T}V_{n+1}$. These assumptions imply that $\sum_{U}V_{n}$ is a strict greatest least lower bound of $\mathcal{B}(U,{\langle}V_{n}\mid n<\omega{\rangle})$. This shows that the cofinal type of $\sum_{U}V_{n}$ is much more complicated than the cofinal type of $U\cdot V$. We also provide an example (Prop. 1.21) where $U<_{T}\sum_{U}V_{n}<U\times\prod_{n\in X}V_{n}$ for every $X\in U$.
Finally, we study the class of ultrafilters $U$ such that $U\geq_{T}\omega^{\omega}$. The partial order $\omega^{\omega}$ appeared quite a bit in the literature [35, 24, 25] in the context of the Tukey order on Borel ideals. In the context of general ultrafilters on $\omega$, it is known to be the immediate successor of the Tukey type $\omega$ [24]
Theorem 0.1 (Louveau-Velickovic).
If $I$ is any ideal such that $I<_{T}\omega^{\omega}$ then $I$ is countably generated.
this partial order appeared in the work of Molvice [28, Question 4.7], and Dobrinen and Todorcevic who proved the following [14, Thm. 35]:
Theorem 0.2 (Dobrinen-Todorcevic).
The following are equivalent for a $p$-point:
  1. (1)
    $U\cdot U\equiv_{T}U$.
  2. (2)
    $U\geq_{T}\omega^{\omega}$.
They then proved that rapid ultrafilters are above $\omega^{\omega}$ and deduced that rapid $p$-points satisfy $U\cdot U\equiv_{T}U$. This was lately improved by Benhamou and Dobrinenn [3, Thm 1.18] to the general setup of $I$-p.i.p.
Theorem 0.3 (Dobrinen-B.).
Let $U$ be an ultrafilter. Then the following are equivalent:
  1. (1)
    $U\cdot U\equiv_{T}U$.
  2. (2)
    There is an ideal $I\subseteq U^{*}$ such that $U\geq_{T}I^{\omega}$ and $U$ has the $I$-p.i.p.
Taking $I=\mathrm{fin}$ reproduces the difficult part from Dobrinen and Todorcevic’s result [^6] Indeed, as pointed out by Dobrinen and Todorcevic, it is easy to see that $U\cdot U\geq_{T}\omega^{\omega}$, so $(1)\Rightarrow(2)$ is straightforward.. In  §4, we first note that in some sense the assumption that $U$ is $I$-p.i.p in the theorem above is not optimal.
Theorem.
Let $U$ be any ultrafilter over a set $X$. If $U$ does not have the $I$-p.i.p then $U\geq_{T}\prod_{x\in X}I$.
In particular, if $U$ is not a $p$-point then automatically $U\geq_{T}\omega^{\omega}$. Hence, to understand the class of ultrafilters Tukey above $\omega^{\omega}$, it suffices to restrict our attention to $p$-points. Inside the class of $p$-point, it is known to be consistent that there are $p$-point which are not above $\omega^{\omega}$ (see [14]), and the class of rapid p-point is currently the largest subclass of $p$-points known to be above $\omega^{\omega}$. In the second part of Section 4, we enlarge this class by introducing the notion of $\alpha$-almost-rapid (Definition 4.11), which is a weakening of rapidness.
Theorem.
Suppose that $U$ is $\alpha$-almost-rapid, then $U\geq_{T}\omega^{\omega}$.
Finally, we prove that the class of almost rapid ultrafilters is a strict extension of the class of rapid ultrafilters, even among $p$-points.
Theorem.
Assume CH. Then there is a non-rapid almost-rapid $p$-point ultrafilter.
This theorem produces a large class of ultrafilters which are above $\omega^{\omega}$. This paper is organized as follows:
  • In §1, we start with some preliminary definitions and known results. The main goal of this section is the investigation of the cofinal type of $\sum_{U}V_{\alpha}$.
  • In §2, we provide a systematic study of the $I$-p.i.p and deterministic ideals.
  • In §3, we prove our main results.
  • In §4 we investigate the class of ultrafilters Tukey above $\omega^{\omega}$.
  • In §5 we present some open problems and possible directions.
Notations
$[X]^{<\lambda}$ denotes the set of all subsets of $X$ of cardinality less than $\lambda$. Let $\mathrm{fin}=[\omega]^{<\omega}$, and $\mathrm{FIN}=\mathrm{fin}\setminus\{\emptyset\}$. For a collection of sets $(P_{i})_{i\in I}$ we let $\prod_{i\in I}P_{i}=\{f:I\rightarrow\bigcup_{i\in I}P_{i}\mid\forall i,\,f(i)\in P_{i}\}$. If $P_{i}=P$ for every $i$, then $P^{I}=\prod_{i\in I}P$. Given a set $X\subseteq\omega$, such that $|X|=\alpha\leq\omega$, we denote by ${\langle}X(\beta)\mid\beta<\alpha{\rangle}$ be the increasing enumeration of $X$. Given a function $f:A\rightarrow B$, for $X\subseteq A$ we let $f^{\prime\prime}X=\{f(x)\mid x\in X\}$, for $Y\subseteq B$ we let $f^{-1}Y=\{x\in X\mid f(x)\in Y\}$, and let $\operatorname{rng}(f)=f^{\prime\prime}A$. Given sets $\{A_{i}\mid i\in I\}$ we denote by $\biguplus_{i\in I}A_{i}$ the union of the $A_{i}$’s when the sets $A_{i}$ are pairwise disjoint. Two partially ordered set $\mathbb{P},\mathbb{Q}$ are isomorphic, denoted by $\mathbb{P}\simeq\mathbb{Q}$, if there is a bijection $f:\mathbb{P}\rightarrow\mathbb{Q}$ which is order-preserving.
1. On the Cofinal types of Fubini sums of ultrafilters
1.1. Some basic definitions and facts
The principal operation we are considering in this paper is the Fubini/tensor sums and products of ultrafilters.
Definition 1.1.
Suppose that $F$ is a filter over an infinite set $X$ and for each $x\in X$, $G_{x}$ is a filter over an infinite set $Y_{x}$. We denote by $\sum_{F}G_{x}$ the filter over $\bigcup_{x\in X}\{x\}\times Y_{x}$, defined by
$A\in\sum_{F}G_{x}\text{ if and only if }\{x\in X\mid(A)_{x}\in G_{x}\}\in F$
[…]
We distinguish here between $F\cdot G$ and $F\times G$ which is the cartesian product of $F$ and $G$ with the order defined pointwise [^7] There are papers which consider the filter $\{A\times B\mid A\in F,B\in G\}$ and denote it by $F\times G$, this filter will not be considered in this paper so there is no risk of confusion..
A filter $F$ on a regular cardinal $\kappa\geq\omega$ is called uniform [^8] Or non-principal. if $J^{*}_{bd}=\{X\subseteq\kappa\mid\kappa\setminus X\text{ is bounded in }\kappa\}\subseteq U$.
Definition 1.2.
Let $F$ be a filter over a regular cardinal $\kappa\geq\omega$.
  1. (1)
    $F$ is $\lambda$-complete if $F$ is closed under intersections of less than $\lambda$ many of its members.
  2. (2)
    $F$ is Ramsey if for any function $f:[\kappa]^{2}\rightarrow 2$ there is an $X\in F$ such that $f\restriction[X]^{2}$ is constant.
  3. (3)
    $F$ is selective if for every function $f:\kappa\rightarrow\kappa$, there is an $X\in F$ such that $f\restriction X$ is either constant or one-to-one.
  4. (4)
    (Kanamori [21]) $F$ is rapid if for each normal function $f:\kappa\rightarrow\kappa$ (i.e. increasing and continuous), there exists an $X\in F$ such that $\text{otp}(X\cap f(\alpha))\leq\alpha$ for each $\alpha<\kappa$. (i.e. bounded pre-images), there is an $X\in F$ such that $|f^{-1}(\{\alpha\})\cap X|\leq\alpha$ for every $\alpha<\kappa$.
  5. (5)
    $F$ is a $p$-point if whenever $f:\kappa\rightarrow\kappa$ is unbounded [^9] Namely, $f^{-1}[\alpha]\notin F$ for every $\alpha<\kappa$ on a set in $F$, it is almost one-to-one mod $F$, i.e. there is an $X\in F$ such that for every $\gamma<\kappa$, $|f^{-1}[\gamma]\cap X|<\kappa$.
  6. (6)
    $U$ is a $q$-point if every function $f:\kappa\rightarrow\kappa$ which is almost one-to-one mod $F$ is injective mod $F$.
a $\kappa$-filter is a uniform, $\kappa$-complete filter.
The following facts are well known.
Fact 1.3.
The following are equivalent for a $\kappa$-ultrafilter $U$:
  1. (1)
    $U$ is Ramsey.
  2. (2)
    $U$ is selective.
  3. (3)
    $U$ is a $p$-point and a $q$-point.
Fact 1.4.
Suppose that $U,V_{\alpha}$ are ultrafilters on $\kappa\geq\omega$ where each $V_{\alpha}$ is uniform. Then $\sum_{U}V_{\alpha}$ is not a $p$-point.
Indeed the function $\pi_{1}$, the projection to the first coordinate, is never almost one-to-one on a set in $X\in\sum_{U}V_{\alpha}$.
Definition 1.5.
Let $F,G$ be filters on $X,Y$ resp. We say that $F$ is Rudin-Keisler below $G$, denoted by $F\leq_{RK}G$, if there is a Rudin-Keisler projection $f:Y\rightarrow X$ such that
$f_{*}(G):=\{A\subseteq X\mid f^{-1}[A]\in G\}=F$
We say that are RK-isomorphic, and denote it by $F\equiv_{RK}G$ if there is a bijection $f$ such that $f_{*}(F)=G$.
It is well known that if $F\leq_{RK}G\wedge G\leq_{RK}F$ then $F\equiv_{RK}G$ and that $F,G\leq_{RK}F\cdot G$ via the projection to the first and second coordinates respectively. Also, the Rudin-Keisler order implies the Tukey order. A Ramsey ultrafilter over $\kappa$ is characterized as being Rudin-Keisler minimal among $\kappa$-ultrafilters.
Next, let us record some basic terminology and facts regarding cofinal types. Given two directed partially ordered sets $\mathbb{P},\mathbb{Q}$, the Cartesian product $\mathbb{P}\times\mathbb{Q}$ ordered pointwise, is the least upper bound of $\mathbb{P},\mathbb{Q}$ in the Tukey order (see [13]). It follows that $F\times G\leq_{T}F\cdot G$. More generally, for partially ordered sets $(\mathbb{P}_{i},\leq_{i})$ for $i\in I$, we denote by $\prod_{i\in I}(\mathbb{P}_{i},\leq_{i})$ to be the order over the underlining set $\prod_{i\in I}\mathbb{P}_{i}$ with the everywhere domination order, namely $f\leq g$ iff for all $i\in I$, $f(i)\leq_{i}g(i)$. If the order is clear from the context we omit it and just write $\prod_{i\in I}\mathbb{P}_{i}$. This is the case when $\mathbb{P}_{i}=U_{i}$ is a filter ordered by reversed inclusion of an ideal ordered by inclusion. If for every $i\in I$, $\mathbb{P}_{i}=\mathbb{P}$ we simply write $\mathbb{P}^{I}$.
1.2. The cofinal type of sums and products
The following theorem [14] provides the starting point for the analyses of the cofinal type of sums of ultrafilters:
Theorem 1.6 (Dobrinen-Todorcevic).
Let $F,G_{x}$ be filters as in Definition 1.1. Then:
  1. (1)
    $\sum_{F}G_{x}\leq_{T}F\times\prod_{x\in X}G_{x}$.
  2. (2)
    If $G_{x}=G$ for every $x$, then $F\cdot G\leq_{T}F\times G^{X}$.
  3. (3)
    If $F=G$, then $F\cdot F\leq_{T}F^{X}$.
These Tukey types are invariant under Rudin-Keisler isomorphic copies of the ultrafilters involved, hence we may assume for the rest of this paper that ultrafilters are defined on regular (infinite) cardinals. It was later discovered [29] that $(2),(3)$ of Theorem 1.6 are in fact an equivalences:
Theorem 1.7 (Milovich).
Let $F,G$ are $\kappa$-filters, then $F\cdot G\equiv_{T}F\times G^{\kappa}$ and in particular $F\cdot F\equiv_{T}F^{\kappa}$.
The proof of Milovich’s Theorem go through in case $F$ is any ultrafilter over $\lambda$ and $G$ is $\lambda$-complete.
Corollary 1.8 (Milovich).
For any two $\kappa$-filters $F,G$, $F\cdot(G\cdot G)\equiv_{T}F\cdot G$ and $F^{3}\equiv_{T}F^{2}$.
It is tempting to conjecture that $\sum_{U}V_{\alpha}\equiv_{T}U\times\prod_{\alpha<\lambda}V_{\alpha}$, however, this will not be the case in general, as indicated by the following example:
Example 1.9.
Suppose that $U$ and $V$ are Tukey incomparable ultrafilters on $\omega$, and $U\equiv_{T}U\cdot U$. This situation is obtained for example under $Cov(\mathcal{M})=\mathfrak{c}$  [^10] By Ketonen [22], this assumption implies that there are $(2^{c})^{+}$-many distinct selective ultrafilters. Then there are two Tukey incomparable selective ultrafilters and by Dobrinen and Todorcevic [14], $U\cdot U\equiv_{T}U$ for any selective ultrafilter.. The incomparability requirement ensures that $U\times V>_{T}U$. Let $V_{0}=V$ and $V_{n}=U$ for $n>0$. Then
$\sum_{U}V_{n}=U\cdot U\equiv_{T}U<_{T}U\times V\leq_{T}U\times\prod_{n<\omega}V_{n}.$
The point of the example is that the sum is insensitive to removing a neglectable set of coordinates, while the product changes if we remove even a single coordinate. Another quite important difference is that $U\times\prod_{x\in X}V_{x}$ is insensitive to permutations of the indexing set, while $\sum_{U}V_{x}$ is. Formally, this is expressed by the following fact:
Fact 1.10.
Let $U$ be an ultrafilter over $\lambda\geq\omega$ and $U_{\alpha}$ on $\delta_{\alpha}$. For every $X\in U$, $\sum_{U}V_{\alpha}\leq_{T}U\times\prod_{\alpha\in X}V_{\alpha}\leq_{T}U\times\prod_{\alpha<\lambda}V_{\alpha}$.
Proof.
The right inequality is clear. The left one is also simple, since the set $\mathcal{X}\subseteq\sum_{U}V_{\alpha}$, of all $Y$ such that $\pi_{1}^{\prime\prime}Y\subseteq X$ is a cofinal set in $\sum_{U}V_{\alpha}$ and therefore the map $F:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow\sum_{U}V_{\alpha}$ defined by
$F({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})=\bigcup_{\alpha\in Z\cap X}\{\alpha\}\times A_{\alpha}$
is monotone and has cofinal image. ∎
In this section, we provide further insight into the cofinal type of $\sum_{U}V_{\alpha}$. We will focus on $\kappa$-ultrafilters, so our initial assumption is that $U$ is a $\lambda$-ultrafilter for $\lambda\geq\omega$ and ${\langle}V_{\alpha}\mid\alpha<\lambda{\rangle}$ is a sequence of ultrafilters such that each $V_{\alpha}$ is a $\delta_{\alpha}$-ultrafilter where $\delta_{\alpha}\geq\omega$. Towards our first result, consider the set
$\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})=\{U\times\prod_{\alpha\in X}V_{\alpha}\mid X\in U\}$
ordered by the Tukey order. This is clearly a downward-directed set. Our goal is to prove that in some sense, $\sum_{U}V_{\alpha}$ is the greatest lower bound of $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$. Consider the maps
$\pi_{X}:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow\sum_{U}V_{\alpha},\ \ \ \ \pi_{X,Y}:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow U\times\prod_{\alpha\in Y}V_{\alpha}$
Defined for $X,Y\in U$ where $Y\subseteq X$ defined by
$\pi_{X}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})=\bigcup_{\alpha\in X\cap Z}\{\alpha\}\times A_{\alpha}\ \text{ and }$
$\pi_{X,Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})={\langle}Z,{\langle}A_{\alpha}\mid\alpha\in Y{\rangle}{\rangle}.$
Then
  1. (1)
    $\pi_{X}$ is monotone cofinal and $\operatorname{rng}(\pi_{X})$ is exactly all the sets $B\in\sum_{U}V_{\alpha}$ in standard form [^11] A set $B\in\sum_{U}V_{\alpha}$ is said to be in standard form if for every $\alpha<\lambda$, either $(B)_{\alpha}=\emptyset$ or $(B)_{\alpha}\in V_{\alpha}$. such that $\pi^{\prime\prime}B\subseteq X$.
  2. (2)
    $\pi_{X,Y}$ is monotone cofinal.
  3. (3)
    $\pi_{Y}\circ\pi_{X,Y}(C)\subseteq\pi_{X}(C)$.
[…]
$f_{Y}(\pi_{X,Y}(C))=g(\pi_{Y}(\pi_{X,Y}(C))\geq_{\mathbb{P}}g(\pi_{X}(C))=f_{X}(C)$
Definition 1.11.
A sequence of monotone cofinal maps
${\langle}f_{X}:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow\mathbb{P}\mid X\in U{\rangle}$
if said to be coherent if
$({\dagger})\ \ \text{whenever }Y\subseteq X,\text{ and }C\in U\times\prod_{\alpha\in X}V_{\alpha},\ f_{Y}(\pi_{X,Y}(C))\geq_{\mathbb{P}}f_{X}(C).$
A poset $\mathbb{P}$ is said to be uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if there is a coherent sequence of monotone cofinal maps ${\langle}f_{X}:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow\mathbb{P}\mid X\in U{\rangle}$.
The following theorem says that $\sum_{U}V_{\alpha}$ is the greatest lower bound among all the posts uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$.
Theorem 1.12.
Suppose that $\mathbb{P}$ is a complete order. Then $\mathbb{P}$ is uniformly bellow $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
Proof.
From right to left was already proven in the paragraph before Definition 1.11. Let us prove from left to right. Let ${\langle}f_{X}\mid X\in U{\rangle}$ be the sequence witnessing that $\mathbb{P}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$. Let $\mathcal{X}\subseteq\sum_{U}V_{\alpha}$ be the usual cofinal set of all the sets $A\in\sum_{U}V_{\alpha}$ is a standard form. Let us define $F:\chi\rightarrow\mathbb{P}$ monotone and cofinal,
$F(A)=f_{\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}(A)_{\alpha}\mid\alpha\in\pi^{\prime\prime}A{\rangle}{\rangle})$
We claim first (and most importantly) that $F$ is monotone. Suppose that $A,B\in\chi$ are such that $A\subseteq B$. Then,
  1. (a.)
    $\pi_{1}^{\prime\prime}A\subseteq\pi_{1}^{\prime\prime}B$ and
  2. (b.)
    for every $\alpha<\lambda$, $(A)_{\alpha}\subseteq(B)_{\alpha}$.
Define the sequence ${\langle}X_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}$ by $X_{\alpha}=(A)_{\alpha}$ for $\alpha\in\pi_{1}^{\prime\prime}A$ and $X_{\alpha}=(B)_{\alpha}$ for $\alpha\in\pi_{1}^{\prime\prime}B\setminus\pi_{1}^{\prime\prime}A$. Note that
$\pi_{\pi_{1}^{\prime\prime}B,\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}X_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}{\rangle})={\langle}\pi_{1}^{\prime\prime}A,{\langle}(A)_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}A{\rangle}{\rangle}$
[…]
$F(A)=f_{\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}(A)_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}A{\rangle}{\rangle})=f_{\pi_{1}^{\prime\prime}A}(\pi_{\pi_{1}^{\prime\prime}B,\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}X_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}{\rangle})$
$\geq_{\mathbb{P}}f_{\pi_{1}^{\prime\prime}B}({\langle}\pi_{1}^{\prime\prime}A,{\langle}X_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}{\rangle})\geq_{\mathbb{P}}f_{\pi_{1}^{\prime\prime}B}({\langle}\pi_{1}^{\prime\prime}B,{\langle}(B)_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}{\rangle}=F(B)$
To see it is cofinal, let $p\in\mathbb{P}$ be any element, fix any $X\in U$, since $f_{X}$ is cofinal, there is ${\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}\in U\times\prod_{\alpha\in X}V_{\alpha}$ such that $f_{X}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})\geq_{\mathbb{P}}p$. Consider $A=\cup_{\alpha\in Z\cap X}\{\alpha\}\times A_{\alpha}$. Then
$F(A)=f_{Z\cap X}({\langle}Z\cap X,{\langle}A_{\alpha}\mid\alpha\in Z\cap X{\rangle}{\rangle})=$
$f_{Z\cap X}(\pi_{X,Z^{\prime}\cap X}({\langle}Z\cap X,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}))\geq_{\mathbb{P}}f_{X}({\langle}Z\cap X,{\langle}A_{\alpha}\mid\alpha\in X{\rangle})\geq_{\mathbb{P}}p$
Lemma 1.13.
Suppose that $\mathbb{P}$ is complete and for each $X\in U$, $\mathcal{X}_{X}\subseteq U\times\prod_{\alpha\in X}V_{\alpha}$ is such that:
  1. (1)
    $\mathcal{X}_{X}$ is a cofinal subset of $U\times\prod_{\alpha\in X}V_{\alpha}$.
  2. (2)
    $f_{X}:\mathcal{X}_{X}\rightarrow\mathbb{P}$ is monotone cofinal.
  3. (3)
    whenever $Y\subseteq X$, $\pi_{X,Y}^{\prime\prime}\mathcal{X}_{X}\subseteq\mathcal{X}_{Y}$ and $f_{Y}(\pi_{X,Y}(C))\geq_{\mathbb{P}}f_{X}(C)$
Then $\mathbb{P}$ is uniformly below $\mathcal{B}(U,{\langle}B_{\alpha}\mid\alpha<\lambda{\rangle})$.
Proof.
Define $f^{*}_{X}:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow\mathbb{P}$ by
$f^{*}_{X}(A)=\sup\{f_{X}(B)\mid A\subseteq B\in\mathcal{X}_{X}\}.$
[…]
$f^{*}_{X}(C)=\sup\{f_{X}(B)\mid C\subseteq B\in\mathcal{X}_{X}\}\leq\sup\{f_{Y}(\pi_{X,Y}(B))\mid C\leq_{\mathbb{P}}B\in\mathcal{X}_{X}\}$
$\leq_{\mathbb{P}}\sup\{f_{Y}(B^{\prime})\mid\pi_{X,Y}(C)\subseteq B^{\prime}\in\mathcal{X}_{Y}\}=f^{*}_{Y}(\pi_{X,Y}(C)).$
Hence ${\langle}f_{X}^{*}\mid X\in U{\rangle}$ is coherent and therefore $\mathbb{P}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$. ∎
Corollary 1.14.
Let $U$ be an ultrafilter on $\lambda\geq\omega$ and that each $V_{\alpha}$ is a $\delta_{\alpha}$-complete ultrafilter on some $\delta_{\alpha}>\alpha$. $\mathbb{P}\leq_{T}V_{\alpha}$ for every $\alpha<\lambda$, then $\mathbb{P}^{\lambda}\leq_{T}\sum_{U}V_{\alpha}$.
Proof.
We fix for every $\alpha<\lambda$, $f_{\alpha}:V_{\alpha}\rightarrow\mathbb{P}$ monotone and cofinal. Now for every $X\in U$, we define a cofinal set $\mathcal{X}_{X}\subseteq U\times\prod_{\alpha\in X}V_{\alpha}$ consisting of all the elements ${\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}\in U\times\prod_{\alpha\in X}V_{\alpha}$ such that for every $\alpha<\beta$ in $X$, $f_{\alpha}(A_{\alpha})\leq_{\mathbb{P}}f_{\beta}(A_{\beta})$.
Claim 1.15.
$\mathcal{X}_{X}$ is cofinal in $U\times\prod_{\alpha\in X}V_{\alpha}$
Proof.
Let ${\langle}Z,{\langle}B_{\alpha}\mid\alpha\in X{\rangle}{\rangle}$, let us construct $A_{\alpha}$ recursively. Let $B_{0}=A_{0}$. Suppose that $A_{\alpha}$ for $\alpha\in X\cap\beta$ where defined for some $\beta\in X$. Then for each $\alpha\in X\cap\beta$, we find (by cofinality of $f_{\beta}$) a set $C_{\alpha}\in V_{\beta}$ such that $f_{\alpha}(A_{\alpha})\leq_{\mathbb{P}}f_{\beta}(C_{\alpha})$. Let $A_{\beta}=B_{\beta}\cap\bigcap_{\alpha<\beta}C_{\alpha}$. By $\delta_{\beta}$-completeness of $V_{\beta}$, $A_{\beta}\in V_{\beta}$. By monotonicity of $f_{\beta}$, we conclude that for every $\alpha<\beta$, $f_{\alpha}(A_{\alpha})\leq_{\mathcal{P}}f_{\beta}(A_{\beta})$. It is now clear that ${\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}\in\mathcal{X}_{X}$ and above ${\langle}Z,{\langle}B_{\alpha}\mid\alpha\in X{\rangle}{\rangle}$. ∎
Note that $\pi_{X,Y}^{\prime\prime}\mathcal{X}_{X}\subseteq\mathcal{X}_{Y}$. Define $f_{X}:\mathcal{X}_{X}\rightarrow\mathbb{P}^{\lambda}$ by
$f_{X}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})={\langle}f_{X(\alpha)}(A_{X(\alpha)})\mid\alpha<\lambda{\rangle}.$
Let ${\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}\in\mathcal{X}_{X}$, and let $Y\subseteq X$, then $Y(\alpha)\geq X(\alpha)$. Hence, by definition of $\mathcal{X}_{X}$, $f_{X(\alpha)}(A_{X(\alpha)})\leq_{\mathbb{P}}f_{Y(\alpha)}(A_{Y(\alpha)})$. We conclude that
$f_{X}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})={\langle}f_{X(\alpha)}(A_{X(\alpha)})\mid\alpha<\lambda{\rangle}\leq_{\mathbb{P}^{\omega}}{\langle}f_{Y(\alpha)}(A_{Y(\alpha)})\mid\alpha<\lambda{\rangle}$
$=f_{Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in Y{\rangle}{\rangle})=f_{Y}(\pi_{X,Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}))$
Hence by Lemma 1.13 $\mathbb{P}^{\lambda}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ and by Theorem 1.12, $\mathbb{P}^{\lambda}\leq_{T}\sum_{U}V_{\alpha}$. ∎
In particular, $U_{\alpha}$ is Tukey-top for a set of $\alpha$’s in $U$, then $\sum_{U}U_{\alpha}$ is Tukey top.
It is unclear whether being uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ is equivalent to simply being a Tukey below each $X\in\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$. Hence it is unclear if $\sum_{U}V_{\alpha}$ is indeed the greatest lower bound of $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ in the usual sense; if every $\mathbb{P}$ which is a lower bound in the Tukey order for $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ is Tukey below $\sum_{U}V_{\alpha}$. Let us give a few common configurations of the Tukey relation among the ultrafilters $V_{\alpha}$ in which $\sum_{U}V_{\alpha}$ is the greatest lower bound in the usual sense. Let us denote that by $\sum_{U}V_{\alpha}=\inf(\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle}))$.
The following is a straightforward corollary from Theorem 1.12:
Corollary 1.16.
Let $X_{0}\in U$, then $\sum_{U}V_{\alpha}\equiv_{T}U\times\prod_{\alpha\in X_{0}}V_{\alpha}$ if and only if $U\times\prod_{\alpha\in X_{0}}V_{\alpha}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$. In that case $\sum_{U}V_{\alpha}=\inf(\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$.
The second case in which $\sum_{U}V_{\alpha}$ turns out to be the greatest lower bound is the following:
Lemma 1.17.
Suppose that there is a set $X_{0}\in U$ such that for every $\alpha<\beta\in X_{0}$, $V_{\alpha}$ is a $\kappa$-complete ultrafilter such that $V_{\alpha}\cdot V_{\alpha}\equiv_{T}V_{\alpha}>_{T}V_{\beta}$. Then $\sum_{U}V_{\alpha}=\inf(\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle}))$ is a strict greatest lower bound.
Proof.
First note that for every $Y\subseteq X$, by the assumptions,
$V_{\min(Y)}\leq_{T}\prod_{m\in Y}V_{m}\leq_{T}\prod_{m\in Y}V_{\min(Y)}\equiv_{T}V_{\min(Y)}\cdot V_{\min(Y)}\equiv_{T}V_{\min(Y)}.$
Therefore, if $\mathbb{P}\leq_{T}B$ for every $B\in\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ then $\mathbb{P}\leq_{T}V_{\alpha}$ for every $\alpha\in X$. By corollary 1.14, it follows than that $\mathbb{P}\leq_{T}\sum_{U}V_{\alpha}$. Moreover, $\sum_{U}V_{\alpha}$ is strictly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$, since for every $\beta<\lambda$, $\sum_{U}V_{\alpha}\leq_{T}V_{\beta+1}<_{T}V_{\beta}$. ∎
We will later show that the assumptions in the above Lemma are consistent. Before that, we consider the third configuration in which the ultrafilters are increasing, the proof below works only for ultrafilters on $\omega$ and we do not know whether it is possible to generalize it to other ultrafilters.
Lemma 1.18.
Suppose that $U,V_{n}$ are ultrafilters on $\omega$, such that on a set $X_{0}\in U$, for every $n\leq m\in X_{0}$, $V_{n}\leq_{T}V_{m}$. Then
$U\times\prod_{n\in X_{0}}V_{n}\equiv_{T}\sum_{U}V_{n}=\inf(\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle}))$
Proof.
By Corollary 1.16, if $\sum_{U}V_{n}\equiv_{T}U\times\prod_{n\in X_{0}}V_{\alpha}$, then it must be the greatest lower bound as well. To prove the Tukey-equivalence, first note that $\sum_{U}V_{n}\leq_{T}U\times\prod_{n\in X_{0}}V_{n}$ by Fact 1.10. For the other direction, define for every $n\in X_{0}$, $n^{+}=\min(X_{0}\setminus n+1)$ and let $f_{n^{+},n}:V_{n^{+}}\rightarrow V_{n}$ monotone and cofinal. Denote by $n^{+2}=(n^{+})^{+}$ and $n^{+k}=(n^{+(k-1)})^{+}$ be the $k^{\text{the}}$ successor of $n$ in $X_{0}$. For any $n<m\in X_{0}$, suppose that $m=n^{+k}$ and let $f_{m,n}=f_{n^{+k},n^{+(k-1)}}\circ f_{n^{+(k-1)},n^{+(k-1)}}\circ...\circ f_{n^{+},n}$. Moreover, let $f_{n,n}:V_{n}\rightarrow V_{n}$ be the identity. Hence $f_{m,n}:V_{m}\rightarrow V_{n}$ is monotone cofinal, and if $k\in X_{0}\cap[n,m]$ then $f_{m,n}=f_{k,n}\circ f_{m,k}$.
Let us define a coherent sequence of cofinal maps from a cofinal subset of $U\times\prod_{n\in X}V_{n}$ to $U\times\prod_{n\in X_{0}}V_{n}$ for $X\in U$. Consider the collection $\mathcal{X}_{X}\subseteq U\times\prod_{n\in X}V_{n}$ of all ${\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}$ such that for all $n,m\in X\cap X_{0}$, if $n<m$ then $f_{m,n}((A)_{m})\subseteq(A)_{n}$. It is straightforward to check that if $Y\subseteq$ then $\pi_{X,Y}^{\prime\prime}\mathcal{X}_{X}\subseteq\mathcal{X}_{Y}$.
Claim 1.19.
$\mathcal{X}_{X}$ is cofinal in $U\times\prod_{n\in X}V_{n}$.
Proof.
Let ${\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}\in U\times\prod_{n\in X}V_{n}$. We define a sequence $X_{n}$ be induction on $n\in X$. $X_{n_{0}}=(A)_{n_{0}}$. Suppose we have defined $X_{n_{k}}\in V_{n_{k}}$ for some $k<m$, For each $k$, we find $C_{m,k}\in V_{m}$ such that $f_{m,k}(C_{m,k})\subseteq X_{n_{k}}$. Define $X_{n_{m}}=(A)_{n_{m}}\cap(\bigcap_{k<m}C_{m,k})$. By monotonicity of the $f_{m,k}$’s $f_{m,k}(X_{n_{m}})\subseteq X_{n_{k}}$. Let $A_{1}={\langle}Z,{\langle}X_{n}\mid n\in X{\rangle}{\rangle}$, then $A_{1}\in\mathcal{X}$ and $A_{1}\geq{\langle}Z,{\langle}A_{\alpha}\mid n\in X{\rangle}{\rangle}$. ∎
Fix the unique order isomorphism $\sigma_{X,X_{0}}:X\rightarrow X_{0}$ (which then satisfy $\sigma(n)\leq n$ as $X\subseteq X_{0}$) and let $f_{X}:\mathcal{X}_{X}\rightarrow U\times\prod_{n\in X_{0}}V_{n}$ be defined by
$f_{X}({\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle})={\langle}Z,{\langle}f_{\sigma_{X,X_{0}}^{-1}(n),n}((A)_{\sigma_{X,X_{0}}^{-1}(n)})\mid n\in X_{0}{\rangle}{\rangle}.$
Clearly, $f_{X}$ is monotone, let us check that it is cofinal and that the sequence ${\langle}f_{X}\mid X\in U\restriction X_{0}{\rangle}$ is coherent. Suppose that $C_{1}={\langle}Z,{\langle}A_{n}\mid n\in X_{0}{\rangle}{\rangle}\in U\times\prod_{n\in X_{0}}V_{n}$. , We find ${\langle}B_{n}\mid n\in X_{0}{\rangle}\geq{\langle}A_{n}\mid n\in X_{0}{\rangle}$ such that $f_{m,n}(B_{m})\subseteq B_{n}$ for every $n<m\in X_{0}$. This is possible as before, constructing the $B_{n}$’s by induction and the fact that at each step we only have finitely many requirements, so we can intersect the corresponding finitely many sets. Now take ${\langle}Z,{\langle}B_{n}\mid n\in X{\rangle}{\rangle}\in\mathcal{X}_{X}$. Then
$f_{X}({\langle}Z,{\langle}B_{n}\mid n\in X{\rangle}{\rangle})={\langle}Z,{\langle}f_{\sigma_{X,X_{0}}^{-1}(n),n}(B_{\sigma_{X,X_{0}}^{-1}(n)})\mid n\in X_{0}{\rangle}{\rangle}.$
Since $\sigma_{X,X_{0}}^{-1}(n)\geq n$ for every $n\in X_{0}$, we conclude that
$f_{\sigma_{X,X_{0}}^{-1}(n),n}(B_{\sigma_{X,X_{0}}^{-1}(n)})\subseteq B_{n},$
[…]
$f_{\sigma_{Y,X_{0}}^{-1}(n),n}((A)_{\sigma_{Y,X_{0}}^{-1}(n)})=f_{\sigma_{X,X_{0}}^{-1}(n),n}(f_{\sigma_{Y,X_{0}}^{-1}(n),\sigma_{X,X_{0}}^{-1}(n)}(A_{\sigma_{Y,X_{0}}^{-1}(n)}))\subseteq$
$\subseteq f_{\sigma_{X,X_{0}}^{-1}(n),n}(A_{\sigma_{X,X_{0}}^{-1}(n)}).$
It follows that $f_{Y}(\pi_{X,Y}({\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}))\geq f_{X}({\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle})$
The above Lemma recovers Milovich’s theorem 1.7, taking each $V_{n}=V$ for every $n$.
Our next goal is to prove that the assumptions of Lemma 1.17 are consistent. This example shows that the cofinal type of $\sum_{U}V_{n}$ in general can be quite complicated. To do that, we will need a theorem of Raghavan and Todorcevic from [31] regarding the canonization of cofinal maps from basically generated ultrafilters. The notion of basically generated ultrafilters was introduced by Dobrinen and Todorcevic [14] as an attempt to approximate the class of ultrafilters which are not Tukey-top. Recall that an ultrafilter $U$ is called basically generated if there is a cofinal set $\mathcal{B}\subseteq U$ such that for every sequence ${\langle}b_{n}\mid n<\omega{\rangle}\subseteq\mathcal{B}$ which converges [^14] A sequence ${\langle}A_{n}\mid n<\omega{\rangle}$ of subsets $\omega$ is said to converge to $A$ if for every $n<\omega$ there is $N<\omega$ such that for every $m\geq N$, $A_{m}\cap n=A\cap n$. to an element of $\mathcal{B}$, there is $I\in[\omega]^{\omega}$ such that $\cap_{i\in I}A_{i}\in U$. A $p$-point ultrafilter $U$ is basically generated as witnessed by $\mathcal{B}=U$ ([14, Thm. 14]). Dobrinen and Todorcevic proved that produces and sums of $p$-points must also be basically generated ([14, Thm. 16]).
Theorem 1.20 (Raghavan-Todorcevic).
Let $U$ be a basically generated ultrafilter and $V$ be any ultrafilter such that $V\leq_{T}U$. Then there is $P\subseteq\mathrm{FIN}$ such that:
  1. (1)
    $\forall t,s\in P,\ t\subseteq s\Rightarrow t=s$.
  2. (2)
    $V$ is Rudin-Keisler below $U(P)$, namely, there is $f:P\rightarrow\omega$ such that $V=\{X\subseteq\omega\mid f^{-1}[X]\in U(P)\}$.
  3. (3)
    $U(P)\equiv_{T}V$.
Where is the filter $U(P)=\{A\subseteq P\mid\exists a\in U.[a]^{<\omega}\subseteq A\}$.
The forcing notion $P(\omega)/\mathrm{fin}$ consists of infinite sets, ordered by inclusion up to a finite set. Namely, $X\leq^{*}Y$ if $X\setminus Y$ is finite. In the next proposition, we consider the forcing notion $\mathbb{P}=\prod_{n<\omega}P(\omega)/\mathrm{fin}$, where elements of the product have full support. For more information regarding forcing we refer the reader to [23].
The following items summarize the properties of $\mathbb{P}$ which we will need:
  • $\mathbb{P}$ is $\sigma$-closed, and therefore does not add new subsets of $\omega$, and $\omega_{1}$ is preserved.
  • For each $n$, the projection $\pi_{n}$ of $\mathbb{P}$ to the $n^{\text{th}}$ coordinate is a forcing projection from $\mathbb{P}$ to $P(\omega)/\mathrm{fin}$ [^15] A function from $f:\mathbb{P}\rightarrow\mathbb{Q}$ is called a projection of forcing notions if $f$ is order-preserving, $\operatorname{rng}(f)$ is dense in $\mathbb{Q}$, and for every $p\in\mathbb{P}$ and $q\leq_{\mathbb{Q}}p$, there is $p^{\prime}\leq_{\mathbb{P}}p$ such that $f(p^{\prime})\leq_{\mathbb{Q}}q$..
  • If $G\subseteq\mathbb{P}$ is generic over $V$, then $U_{n}:=\overline{\pi_{n}^{\prime\prime}G}=\{X\in P(\omega)\mid\exists f\in G.f(n)\leq^{*}X\}$ is an ultrafilter over $\omega$ in $V[G]$. Moreover, $U_{n}$ is a generic ultrafilter for $P(\omega)/\mathrm{fin}$.
  • Each $U_{n}$ is a selective ultrafilter
  • $U_{n}\notin V[{\langle}U_{m}\mid m\in\omega\setminus\{n\}{\rangle}]$.
Proposition 1.21.
Let $\mathbb{P}$ be a full support product of $\omega$-copies of $P(\omega)/\mathrm{fin}$. Let $G\subseteq\mathbb{P}$ be generic over $V$. Then in $V[G]$ there is a sequence of ultrafilters $V_{n}$, such that $V_{0}>_{T}V_{1}>_{T}V_{2}...$ and $V_{n}\cdot V_{n}\equiv V_{n}$.
Proof.
For each $n<\omega$, $U_{n}$ is a selective ultrafilter and therefore by Theorem 0.2, $U_{n}\cdot U_{n}\equiv_{T}U_{n}\equiv_{T}(U_{n})^{\omega}$. For every $n<\omega$, define [^16] We thank Gabe Goldberg for pointing out this definition of $V_{n}$.
$V_{n}=\sum_{U_{0}}(U_{n+1}\cdot U_{n+2}\cdot...\cdot U_{n+m})_{0<m<\omega}$
Note that each $U_{n+1}\cdot...\cdot U_{n+m}$ is basically generated as the product of $p$-points. Therefore, $V_{n}$ is also basically generated.
Lemma 1.22.
  1. (1)
    $V_{n}\equiv_{T}U_{0}\times\prod_{n<m<\omega}U_{m}$.
  2. (2)
    $V_{n}\cdot V_{n}\equiv_{T}V_{n}$.
  3. (3)
    $V_{0}>_{T}V_{1}>_{T}V_{2}...$.
Proof of Lemma..
For $(1)$, we note that the ultrafilters $U_{n+1}\cdot...\cdot U_{n+m}$ over which we sum in the definition of $V_{n}$ are increasing in the Tukey order. Hence by Lemma 1.18
$V_{n}\equiv_{T}U_{0}\times\prod_{0<m<\omega}U_{n+1}\cdot...\cdot U_{n+m}$
By Milovich’s Theorem 1.7, and by our assumptions, for each $n,m$
$U_{n+1}\cdot...\cdot U_{n+m}\equiv_{T}U_{n+1}\times U_{n+2}\cdot U_{n+2}\times...\times U_{n+m}\cdot U_{n+m}\equiv_{T}U_{n+1}\times...\times U_{n+m}.$
Hence
$V_{n}\equiv_{T}U_{0}\times\prod_{0<m<\omega}U_{n+1}\times...\times U_{n+m}\equiv_{T}U_{0}\times\prod_{0<m<\omega}(U_{n+m})^{\omega}\equiv_{T}$
$\equiv_{T}U_{0}\times\prod_{0<m<\omega}U_{n+m}\cdot U_{n+m}\equiv_{T}U_{0}\times\prod_{0<m<\omega}U_{n+m}$
Now for $(2)$, we use $(1)$. For each $n<\omega$
$V_{n}\cdot V_{n}\equiv_{T}(V_{n})^{\omega}\equiv_{T}(U_{0}\times\prod_{0<m<\omega}U_{n+m})^{\omega}\equiv_{T}(U_{0})^{\omega}\times\prod_{0<m<\omega}(U_{n+m})^{\omega}\equiv_{T}$
$\equiv_{T}U_{0}\times\prod_{0<m<\omega}U_{n+m}\equiv_{T}V_{n}$
For $(3)$, it follows from $(1)$ that
$V_{0}\geq_{T}V_{1}\geq_{T}V_{2}....$
Suppose toward a contradiction that $V_{n}\equiv_{T}V_{n+1}$ for some $n$. Then $U_{n+1}\leq_{T}V_{n+1}$. Note that
$V_{n+1}\in V[U_{0},{\langle}U_{m}\mid n+1<m<\omega{\rangle}].$
By mutual genericity $U_{n+1}\notin V[U_{0},{\langle}U_{m}\mid n+1<m<\omega{\rangle}]$. Since $V_{n+1}$ is basically generated, Theorem 1.20 implies that there is $P\subseteq FIN$ such that $U_{n+1}\leq_{RK}V_{n}(P)$. Note that since $\mathbb{P}$ is $\sigma$-closed, $P\in V$ and therefore $V_{n}(P)\in V[U_{0},{\langle}U_{m}\mid n+1<m<\omega{\rangle}]$. Also the Rudin-Keisler projection $f$ such that $f_{*}(V_{n}(P))=U_{n+1}$ is in the ground model and therefore $U_{n+1}\in V[U_{0},{\langle}U_{m}\mid n+1<m<\omega{\rangle}]$, contradiction. ∎
It follows that $\sum_{U_{0}}V_{n}=\inf(\mathcal{B}(U_{0},{\langle}V_{n}\mid 0<n<\omega{\rangle}))$ is a strict greatest lower bound. Let us also prove that $U_{0}<_{T}\sum_{U}V_{n}$. We will need the following folklore fact.
Fact 1.23.
Suppose that $\sum_{U}V_{n}=\sum_{U}V_{n}^{\prime}$ then $\{n<\omega\mid V_{n}=V^{\prime}_{n}\}\in U$
Proof.
Just otherwise, $Y=\{n<\omega\mid V_{n}\neq V^{\prime}_{n}\}\in U$, in which case, for every $n\in Y$ take $X_{n}\in V_{n}$ such that $X_{n}^{c}\in V^{\prime}_{n}$. Then $A=\bigcup_{n\in Y}\{n\}\times X_{n}\in\sum_{U}V_{n}$, while $A^{\prime}=\bigcup_{n\in Y}\{n\}\times X_{n}^{c}\in\sum_{U}V_{n}^{\prime}$. However $A\cap A^{\prime}=\emptyset$ which contradicts $\sum_{U}V_{n}=\sum_{U}V^{\prime}_{n}$. ∎
Proposition 1.24.
$U_{0}<_{T}\sum_{U_{0}}V_{n}$
Proof.
Otherwise, there would have been a continuous cofinal map $f:U_{0}\rightarrow\sum_{U_{0}}V_{n}$. Since $U_{0}$ is a selective ultrafilter, by Todorcevic [31], if $V\leq_{T}U_{0}$, then there is $\alpha<\omega_{1}$ such that $V=_{RK}U_{0}^{\alpha}$ for some $\alpha<\omega_{1}$. It follows that $\sum_{U_{0}}V_{n}\equiv_{RK}U_{0}^{\alpha}$ for some $\alpha<\omega_{1}$. If $\alpha>1$, then $U_{0}^{\alpha}=\sum_{U_{0}}U_{0}^{\alpha_{n}}$ for some $\alpha_{n}<\alpha$ (The $\alpha_{n}$’s might be constant). It follows that $Y=\{n<\omega\mid V_{n}=_{RK}U_{0}^{\alpha_{n}}\}\in U_{0}$. Since for any $\beta<\omega_{1}$, $U_{0}^{\beta}\in V[U_{0}]$, for any $0<n\in Y$, we conclude that $V_{n}\in V[U_{0}]$ and in particular $U_{1}\in V[U_{0}]$, contradicting the mutual genericity. If $\alpha=1$, then $U_{0}=_{RK}\sum_{U_{0}}V_{n}$ which then implies that $\sum_{U_{0}}V_{n}$ is a $p$-point, in contradiction to Fact 1.4.
2. Two properties of filter
In this section, we present two properties of filters which play a key role in the proof of our main result. The first is the $I$-p.i.p which was introduced in [3], and the second is a new concept called deterministic ideals. Both of them provide an abstract framework in which one can analyze the connection between the Tukey type of an ultrafilter and ideals related to it. We start this section with a systematic study of the $I$-p.i.p. Many of our results in this section generalize to $\kappa$-filters for $\kappa\geq\omega$. However, we will restrict our attention to ultrafilters on $\omega$, as the main application is the main result regarding the commutativity of Fubini products which is already known for $\kappa$-complete ultrafilters on measurable cardinals [2].
2.1. The pseudo intersection property relative to a set.
Given set $\mathcal{F}\subseteq P(X)$, we denote by *=$\mathcal{F}^{*}=\{X\setminus A\mid A\in\mathcal{F}\}$. When $\mathcal{F}$ is a filter, $\mathcal{F}^{*}$ is an ideal which we call the dual ideal, and when $\mathcal{I}$ is an ideal $\mathcal{I}^{*}$ is a filter which we call the dual filter. Ideals are always considered with the (regular) inclusion order.
Fact 2.1.
For every filter $F$, $(F,\subseteq)\simeq(F^{*},\supseteq)$ and in particular $(F,\subseteq)\equiv_{T}(F^{*},\supseteq)$.
Definition 2.2.
A filter $F$ over a countable set $S$ such that $T\subseteq F^{*}$ ($T$ is any subset), is said to satisfy the $T$-pseudo intersection property ($T$-p.i.p) if for every sequence ${\langle}X_{n}\mid n<\omega{\rangle}\subseteq F$, there is $X\in F$ such that for every $n$, there is $t\in T$ such that $X\setminus X_{n}\subseteq t$.
The proof for these simple facts can be found in [3]:
Fact 2.3.
  1. (1)
    Any filter $F$ has the $F^{*}$-p.i.p.
  2. (2)
    $F$ is a $p$-point iff $F$ has $\mathrm{fin}$-p.i.p
The following facts are also easy to verify:
Fact 2.4.
  1. (1)
    If $T$ is downward closed with respect to inclusion, then $F$ has the $T$-p.i.p if and only if for every sequence ${\langle}X_{n}\mid n<\omega{\rangle}\subseteq F$, there is $X\in F$ such that $X\setminus X_{n}\in T$.
  2. (2)
[…]
Benhamou and Dobrinen proved the following:
Proposition 2.5 ([3]).
Suppose that $F$ is a filter and $I\subseteq F^{*}$ is any ideal such that $F$ has the $I$-p.i.p. Then $F^{\omega}\leq_{T}F\times I^{\omega}$.
Theorem 2.6 ([3]).
Suppose that $U$ is an ultrafilter and $I\subseteq U^{*}$ is an ideal such that:
  1. (1)
    $U$ has the $I$-p.i.p.
  2. (2)
    $I^{\omega}\leq_{T}U$
Then $U\cdot U\equiv_{T}U$.
This theorem is the important direction in the equivalence of Theorem 0.3. This subsection is devoted to a systematic study of this property, which will be used in the proof for our main theorem regarding the commutativity of the cofinal types of Fubini products of ultrafilters. First let us provide an equivalent condition to being $I$-p.i.p, similar to the one we have for $p$-points.
Proposition 2.7.
Let $U$ be any ultrafitler. Then the following are equivalent:
  1. (1)
    $U$ has the $I$-p.i.p
  2. (2)
    For any partition ${\langle}A_{n}\mid n<\omega{\rangle}$ such that for any $n$, $A_{n}\notin U$, there is $A\in U$ such that $A\cap A_{n}\in I$ for every $n<\omega$.
  3. (3)
    Every function $f:\omega\rightarrow\omega$ which is unbounded modulo $U$ is $I$-to-one modulo $U$, i.e. there is $A\in U$ such that for every $n<\omega$, $f^{-1}[n+1]\cap A\in I$.
Proof.
The proof is standard and is just a generalization of the usual characterization of $p$-points.
  1. $\underline{(1)\Rightarrow(2)}$
    Let ${\langle}A_{n}\mid n<\omega{\rangle}$ be a partition such that $A_{n}\notin U$. Let $B_{n}=\omega\setminus A_{n}\in U$ and by the $I$-p.i.p there is $A\in U$ such that $A\setminus B_{n}\in I$. It remains to note that $A\setminus B_{n}=A\cap A_{n}$ to conclude $(2)$.
  2. $\underline{(2)\Rightarrow(3)}$
    Let $f:\omega\rightarrow\omega$ be unbounded modulo $U$. Let $A_{n}=f^{-1}[\{n\}]$, then $A_{n}\notin U$. Apply $(2)$ to the partition ${\langle}A_{n}\mid n<\omega{\rangle}$ to find $A\in U$ such that $A\cap A_{n}\in I$. For any $n<\omega$, $f^{-1}[n+1]\cap A=\cup_{m\leq n}f^{-1}[\{m\}]\cap A\in I$. Hence $f$ is $I$-to-one modulo $U$.
  3. $\underline{(3)\Rightarrow(1)}$
    Let ${\langle}B_{n}\mid n<\omega{\rangle}\subseteq U$, and let us assume without loss of generality that it is $\subseteq$-decreasing and that $\bigcap_{n<\omega}B_{n}=\emptyset$. Define $f(n)=\min\{m\mid n\notin B_{m}\}$. Since $\bigcap_{n<\omega}B_{n}=\emptyset$, $f:\omega\rightarrow\omega$ is a well defined function. Apply $(3)$, to find $A\in U$ such that for every $n<\omega$ $f^{-1}[n+1]\cap A\in I$. Now for each $x\in A\setminus B_{n}$, $f(x)\leq n$ and therefore $x\in f^{-1}[n+1]\cap A$ and therefore $A\setminus B_{n}\subseteq f^{-1}[n+1]\cap A\in I$. It follows that $A\setminus B_{n}\in I$ and that $U$ has the $I$-p.i.p.
Proposition 2.8.
  1. (1)
    If $F$ has $T$-p.i.p and $T\subseteq S$, then $F$ has $S$-p.i.p.
  2. (2)
    Suppose that $T_{1},T_{2}\subseteq F^{*}$ are downwards-closed with respect to inclusion, and $F$ has both the $T_{1}$-p.i.p and the $T_{2}$-p.i.p, then $F$ has the $T_{1}\cap T_{2}$-p.i.p
  3. (3)
    Suppose that $f_{*}(G)=F$ and $G$ has the $T$-p.i.p then $F$ has the $\{f^{\prime\prime}t\mid t\in T\}$-p.i.p.
Proof.
For $(1)$, see [3]. For $(2)$, suppose that $U$ has both the $T_{1}$-p.i.p and the $T_{2}$-p.i.p. and $T_{1},T_{2}$ are downwards closed. Let ${\langle}A_{n}\mid n<\omega{\rangle}$ be a sequence, then there are $A,B\in F$ such that for every $n$, $A\setminus A_{n}\in T_{1}$ and $B\setminus A_{n}\in T_{2}$. It follows that $A\cap B\in F$, fix $n<\omega$, then $A\cap B\setminus A_{n}$ is included in both $A\setminus A_{n}$ and $B\setminus A_{n}$ which implies that $A\cap B\setminus A_{n}\in T_{1}\cap T_{1}$ as both $T_{1},T_{2}$ are downwards closed. Hence $F$ has the $T_{1}\cap T_{2}$-p.i.p.
For $(3)$, let ${\langle}X_{n}\mid n<\omega{\rangle}\subseteq F$, then ${\langle}f^{-1}[X_{n}]\mid n<\omega{\rangle}\subseteq G$. Therefore, there is $Y\in G$ such that for every $n$ there is $t_{n}\in T$ such that $Y\setminus f^{-1}[X_{n}]\subseteq t_{n}$. Let $X=f^{\prime\prime}Y\in F$, we have that
$X\setminus X_{n}\subseteq f^{\prime\prime}[Y\setminus f^{-1}[X_{n}]]\subseteq f^{\prime\prime}t_{n}$
Corollary 2.9.
Let $F$ be any filter. Denoted by $U_{F}\subseteq P(F^{*})$ the set generated by all $T$’s such that $F$ has the $T$-p.i.p, namely,
[…]
Then $U_{F}$ is a filter over $F^{*}$.
$U_{F}$ is almost an ultrafilter:
Proposition 2.10.
Suppose that $X_{1},X_{2}\subseteq P(F^{*})$, and $F$ has the $X_{1}\cup X_{2}$-p.i.p, then either $F$ has the $X_{1}$-p.i.p or $X_{2}$-p.i.p
Proof.
Suppose otherwise that $X_{1},X_{2}$ are downward closed, $F$ has the $X_{1}\cup X_{2}$-p.i.p , but does not have the neither the $X_{1}$-p.i.p nor the $X_{2}$-p.i.p. Then there are sequences ${\langle}A_{n}\mid n<\omega{\rangle}$ and ${\langle}B_{n}\mid n<\omega{\rangle}$ such that for every $A,B$ there are $n_{A},m_{B}$ such that for every $t_{1}\in X_{1}$ and every $t_{2}\in X_{2}$, $A\setminus A_{n_{A}}\not\subseteq t_{1}$ and $B\setminus B_{m_{B}}\not\subseteq t_{2}$. Consider the sequence ${\langle}\bigcap_{k\leq n}A_{k}\cap\bigcap_{k\leq n}B_{k}\mid n<\omega{\rangle}$. Then there is $A\in F$ such that for every $l$ there is $t_{l}\in X_{1}\cup X_{2}$ for which $A\setminus\bigcap_{k\leq l}A_{k}\cap\bigcap_{k\leq l}B_{k}\subseteq t_{l}$. For $A$, there are suitable $n_{A},m_{A}$ as above and fix $N=\max(n_{A},m_{A})$. Without loss of generality, $t_{N}\in X_{1}$, in which case, we have $A\setminus A_{n}\subseteq A\setminus\bigcap_{k\leq N}A_{k}\cap\bigcap_{k\leq M}B_{k}\subseteq t_{N}$, contradicting the choice of $n_{A}$. ∎
Note that we cannot ensure that either $X_{1}$ or $X_{2}$ contains a downward closed subset. Our next result investigates how the $I$-p.i.p is preserved under sums of ideals and ultrafilters.
Let $I$ be an ideal on $X$ and for each $x\in X$ let $J_{x}$ be ideals on $Y_{x}$ (resp.). We define the Fubini sum of the ideals $\sum_{I}J_{x}$ over $\bigcup_{x\in X}\{x\}\times Y_{x}$: For $A\subseteq\bigcup_{x\in X}\{x\}\times Y_{x}$,
$A\in\sum_{I}J_{x}\text{ iff }\{x\in X\mid(A)_{x}\notin J_{x}\}\in I.$
We denote by $I\otimes J=\sum_{I}J$. When $I$ is an ideal on $\omega$, we define transifinitely for $\alpha<\omega_{1}$ $I^{\otimes\alpha}$. $I^{\otimes 1}=I$, at the successor step $I^{\otimes(\alpha+1)}=I^{\otimes\alpha}\times I$. At limit step $\alpha$, we fix some cofinal sequence ${\langle}\alpha_{n}\mid n<\omega{\rangle}$ unbounded in $\alpha$ and define $I^{\alpha}=\sum_{I}I^{\otimes\alpha_{n}}$. The above definition of Fubini sum is nothing but the dual operation of the Fubini sum of filters:
Fact 2.11.
$(\sum_{I}J_{x})^{*}=\sum_{I^{*}}J_{x}^{*}$ and in particular $(I\otimes J)^{*}=I^{*}\cdot J^{*}$.
Proposition 2.12.
Let $F,F_{n}$ be filters over countable sets. Suppose that $I\subseteq F^{*}$ and $J_{n}\subseteq F_{n}^{*}$ are ideals for every $n<\omega$. Then if $F$ has $I$-p.i.p and for every $n<\omega$, $F_{n}$ has $J_{n}$-p.i.p, then $\sum_{F}F_{n}$ has $\sum_{I}J_{n}$-p.i.p.
Proof.
Let ${\langle}A_{n}\mid n<\omega{\rangle}$ be a sequence in $\sum_{F}F_{n}$. For each $n$, let $X_{n}=\{m<\omega\mid(A_{n})_{m}\in F_{m}\}\in F$. We find $X\in F$ such that for every $n<\omega$, $X\setminus X_{n}\in I$. For each $m\in X$, we consider $E_{m}=\{n<\omega\mid m\in X_{n}\}$. If $E_{m}$ is finite, we let $Y_{m}=\bigcap_{n\in E_{m}}(A_{n})_{m}\in F_{m}$ (if $E_{m}$ is empty, we let $Y_{m}=\omega$). Otherwise, we find $Y_{m}\in F_{m}$ such that for all $n\in E_{m}$, $Y_{m}\setminus(A_{n})_{m}\in J_{m}$. Let $A=\bigcup_{m\in X}\{m\}\times Y_{m}$. Then Clearly, $A\in\sum_{F}F_{n}$. Let $n<\omega$ and consider $A\setminus A_{n}$. If
$A\setminus A_{n}=(\bigcup_{x\in X\cap X_{n}}\{x\}\times Y_{x}\setminus(A_{n})_{x})\cup(\bigcup_{x\in X\setminus X_{n}}\{x\}\times Y_{x}).$
If $x\notin X\setminus X_{n}$ then $(A\setminus A_{n})_{x}=Y_{x}\setminus(A_{n})_{x}$. Since $x\in X_{n}$, we have $n\in E_{x}$ and therefore $Y_{x}\setminus(A_{n})_{x}\in J_{x}$. We conclude that
$\{x<\omega\mid(A\setminus A_{n})_{x}\notin J_{n}\}=X\setminus X_{n}\in I.$
Hence $A\setminus A_{n}\in\sum_{I}J_{n}$. ∎
One way to obtain non-trivial sets $T$ for which an ultrafilter $U$ has the $T$-p.i.p is by intersecting $U$ with another ultrafilter:
Theorem 2.13.
Suppose that $U_{1},U_{2},...U_{n}$ are any ultrafilters, then for each $1\leq i\leq n$, $U_{i}$ have the $(U_{1}\cap U_{2}\cap...\cap U_{n})^{*}$-p.i.p. In particular $U_{i}^{\omega}\leq_{T}U_{i}\times(U_{1}\cap U_{2}\cap...\cap U_{n})^{*\omega}$.
Proof.
Fix any $1\leq i\leq n$. Note that $(\bigcap_{j=1}^{n}U_{j})^{*}=\bigcap_{j=1}^{n}(U_{j})^{*}$. Suppose otherwise, then there is a sequence ${\langle}X_{n}\mid n<\omega{\rangle}\subseteq U_{i}$ such that for every $X\in U_{i}$ there is $n<\omega$ such that $X\setminus X_{n}\notin\bigcap_{j=1}^{n}U_{j}^{*}$. Since $X\setminus X_{n}\in U_{i}^{*}$, this means that there is $j\neq i$ such that $X\setminus X_{n}\notin U_{j}^{*}$. Since $U_{j}$ is an ultrafilter, it follows that $X\setminus X_{n}\in U_{j}$ and therefore $X\in U_{j}$. We conclude that $U_{i}\subseteq\bigcup_{j\neq i}U_{j}$. There must be $j\neq i$ such that $U_{i}\subseteq U_{j}$, just otherwise, for each $j\neq i$ find $X_{j}\in U_{i}$ such that $X_{j}\notin U_{j}$. Then $X^{*}=\bigcap_{j\neq i}X_{j}\in U_{i}$. But then there is $j^{\prime}\neq i$ such that $X^{*}\in U_{j^{\prime}}$. It follows that $X_{j^{\prime}}\in U_{j^{\prime}}$, contradicting our choice of $X_{j^{\prime}}$. Since ultrafilters are maximal with respect to inclusion, and $U_{i}\subseteq U_{j}$, we conclude that $U_{i}=U_{j}$. Now this is again a contradiction since for some (any) $X$, $X_{n}$ and $X\setminus X_{n}$ disjoint and both in $U$. ∎
The argument above can be generalized to an infinite set of ultrafilters in some cases. Recall that $U$ is an accumulation point (in the topological space $\beta\omega\setminus\omega$) of a set of ultrafilters $\mathcal{A}\subseteq\beta\omega\setminus\omega$ if and only if $U\subseteq\bigcup\mathcal{A}\setminus\{U\}$.
Proposition 2.14.
Suppose that $U$ is not an accumulation point of $\mathcal{A}\subseteq\beta\omega\setminus\omega$. Then $U$ has the $(\bigcap\mathcal{A})^{*}$-p.i.p. In particular $U^{\omega}\leq_{T}U\times(\bigcap\mathcal{A})^{\omega}$.
Proof.
Otherwise, we get that for some sequence ${\langle}X_{n}\mid n<\omega{\rangle}\subseteq U$, for every $X\in U$, there is $n$ such that $X\setminus X_{n}\in V$ for some $V\in\mathcal{A}$. Since $X\setminus X_{n}\notin U$, it follows $V\neq U$. It follows that $U\subseteq\bigcup\mathcal{A}\setminus\{U\}$, contradiction. ∎
Recall that a sequence ${\langle}U_{n}\mid n<\omega{\rangle}$ of ultrafilters on $\omega$ is called discrete if there are disjoint sets $A_{n}\in U_{n}$. This is just equivalent to being a discrete set in the space $\beta\omega\setminus\omega$. In particular, no point $U_{n}$ is in the closure of the others.
Corollary 2.15.
Suppose that $U$ is not in the closure of the $U_{n}$’s, namely, there is a set $A\in U$ such that for every $n$, $A\notin U_{n}$, then $U$ has the $\bigcap_{n<\omega}U_{n}^{*}$-p.i.p.
Corollary 2.16.
If $U_{n}$ is discrete then each $U_{n}$ has the $\bigcap_{n<\omega}U_{n}^{*}$-p.i.p.
The partition given by the discretizing sets of a discrete sequence of ultrafilters can be used to compute the cofinal type of the filter obtained by intersecting the sequence.
Proposition 2.17.
Suppose that $U_{n}$ is discrete. Then $\bigcap_{n<\omega}U_{n}\equiv_{T}\prod_{n<\omega}U_{n}$.
Proof.
Let $A_{n}$ be a partition of $\omega$ so that $A_{n}\in U_{n}$. Define $f({\langle}B_{n}\mid n<\omega{\rangle})=\bigcup_{n<\omega}B_{n}\cap A_{n}$. It is clearly monotone. If $X\in\bigcap_{n<\omega}U_{n}$, we let $X\cap A_{n}=B_{n}\in U_{n}$ Then $X=X\cap\omega=X\cap(\bigcup_{n<\omega}A_{n})=\bigcup_{n<\omega}X\cap A_{n}=f({\langle}B_{n}\mid n<\omega{\rangle})$. To see it is unbounded, suppose that $f^{\prime\prime}\mathcal{A}$ is bounded by $B\in\bigcap_{n<\omega}U_{n}$, then for every $A\in\mathcal{A}$, $A\cap A_{n}\supseteq B\cap A_{n}$. and therefore ${\langle}B\cap A_{n}\mid<\omega{\rangle}$ would bound $\mathcal{A}$. ∎
Corollary 2.18.
If ${\langle}U_{n}\mid n<\omega{\rangle}$ is a discrete sequence of ultrafilters then for every $X\subseteq Y$, $\bigcap_{n\in X}U_{n}\leq_{T}\bigcap_{m\in Y}U_{m}$
2.2. Simple and deterministic ideals
Definition 2.19.
An ideal $I$ is simple if for every ideal $J$, $I\subseteq J$, $I\leq_{T}J$.
Clearly any ultrafilter is simple, also $\mathrm{fin}$ is simple as $\mathrm{fin}\equiv_{T}\omega\leq_{T}F$ for any filter $F$ which is non-principal. To construct examples of ideals which are not simple, we have the following lemma.
Lemma 2.20.
Suppose that ${\langle}U_{n}\mid n<\omega{\rangle}$ is a sequence of discrete ultrafilters, $X\subseteq\omega$ and $n<\omega$ such that $U_{n}\not\leq_{T}\prod_{m\neq n}U_{m}$, then for every $X\cup\{n\}\subseteq Y$, $\bigcap_{n\in X}U_{n}<_{T}\bigcap_{m\in Y}U_{m}$
Proof.
$\bigcap_{n\in X}U_{n}\leq_{T}\bigcap_{m\in Y}U_{m}$ follows from Corollary 2.18. To see that it is strict, suppose otherwise, and let $t\in Y\setminus X$, then $U_{t}\leq_{T}\prod_{m\in Y}U_{m}$ but by assumption $U_{t}\not\leq_{T}\prod_{n\in X}$ and therefore $\bigcap_{n\in X}U_{n}\not\equiv_{T}\bigcap_{m\in Y}U_{m}$. ∎
Clearly, taking any $X\cup\{n\}\subseteq Y$ in the previous lemma, we get $\bigcap_{m\in Y}U_{m}\subseteq\bigcap_{n\in X}U_{n}$. Hence $\bigcap_{m\in Y}U_{m}$ is not simple.
Remark 2.21.
A sequence ${\langle}U_{n}\mid n<\omega{\rangle}$ of $\omega$-many mutually generic ultrafilters for $P(\omega)/\mathrm{fin}$ would be a discrete sequence satisfying the assumptions of Lemma 2.20. To see this, note that by Lemma 1.22 $U_{0}\times\prod_{1<m<\omega}U_{n}$ is Tukey equivalent to some basically generated ultrafilter in $V[U_{0},{\langle}U_{m}\mid m>1{\rangle}]$. Now we argue as in Lemma 1.22, concluding that $U_{1}\not\leq_{T}U_{0}\times\prod_{1<m<\omega}U_{m}$.
Definition 2.22.
We say that an ideal $I$ is deterministic if there is a cofinal set $\mathcal{B}\subseteq I$ such that for every $\mathcal{A}\subseteq\mathcal{B}$, $\bigcup\mathcal{A}\in I$ or $\bigcup\mathcal{A}\in I^{*}$.
Example 2.23.
We claim that $\mathrm{fin}$ is deterministic. Indeed, let $\mathcal{B}=\omega$. Then clearly, $\mathcal{B}$ is a cofinal in $\mathrm{fin}$. Suppose that $\mathcal{A}\subseteq\omega$ is such that $\bigcup\mathcal{A}\notin fin$, then $\mathcal{A}$ is an unbounded set of natural numbers and therefore $\bigcup\mathcal{A}=\omega\in\mathrm{fin}^{*}$. We will see later that for every $\alpha$, $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
The reason that deterministic ideals are interesting is due to the following proposition:
Proposition 2.24.
If $I$ is deterministic then $I$ is simple.
Proof.
Let $I\subseteq J$ and let $\mathcal{B}\subseteq I$ be the cofinal set witnessing that $I$ is deterministic. Let us prove that the identity function $id:\mathcal{B}\rightarrow J$ is unbounded. Suppose that $\mathcal{A}\subseteq\mathcal{B}$ is unbounded, then $\bigcup\mathcal{A}\notin I$, since otherwise, as $\mathcal{B}$ is cofinal in $I$, there would have been $b\in\mathcal{B}$ bounding $\mathcal{A}$. By definition of deterministic ideals, it follows that $\bigcup\mathcal{A}\in I^{*}$, and since $I\subseteq J$, $I^{*}\subseteq J^{*}$ hence $\bigcup\mathcal{A}\in J^{*}$. We conclude that $\bigcup\mathcal{A}\notin J$, namely, $\mathcal{A}$ is unbounded in $J$. Hence the identity function witnesses that $I\equiv_{T}\mathcal{B}\leq_{T}J$. ∎
Proposition 2.25.
Suppose that $I\subseteq X$ is a deterministic ideal over $X$.
  1. (1)
    If $\pi:X\rightarrow Y$ is injective on a set in $I$. Then
    $\pi_{*}(I):=\{a\mid\pi^{-1}[a]\in I\}$
    is deterministic.
  2. (2)
    If $A\subseteq X$, then $I\cap P(A)$ is deterministic.
Proof.
For $(1)$, let $\mathcal{B}\subseteq I$ be a witnessing cofinal set. Let $\mathcal{C}=\{(Y\setminus(\pi^{\prime\prime}[X\setminus b]))\mid b\in\mathcal{B}\}$. Then $\mathcal{C}$ is a cofinal set in $\pi_{*}(I)$. Indeed, if $A\in\pi_{*}(I)$, then $\pi^{-1}[A]\in I$ and there is $b\in\mathcal{B}$ such that $b\supseteq\pi^{-1}[A]$. If $y\notin Y\setminus(\pi^{\prime\prime}X\setminus b)$, then $y=\pi(x)$ for some $x\in X\setminus b$. Since $\pi^{-1}[A]\subseteq b$, then $x\notin\pi^{-1}[A]$ which then implies that $y=\pi(x)\notin A$. We conclude that $A\subseteq Y\setminus(\pi^{\prime\prime}[X\setminus b])\in\mathcal{C}$, as wanted. We claim that $\mathcal{C}$ witnesses that $\pi_{*}(I)$ is deterministic. Let $\mathcal{A}\subseteq\mathcal{B}$ be such that $\bigcup_{a\in\mathcal{A}}Y\setminus(\pi^{\prime\prime}[X\setminus a])\notin\pi_{*}(I)$. Then $\pi^{-1}[\bigcup_{a\in\mathcal{A}}Y\setminus(\pi^{\prime\prime}[X\setminus a])]\notin I$. Simplifying the above set we have
$\pi^{-1}[\bigcup_{a\in\mathcal{A}}Y\setminus(\pi^{\prime\prime}[X\setminus a])]=\pi^{-1}[Y\setminus(\bigcap_{a\in\mathcal{A}}(\pi^{\prime\prime}[X\setminus a]))]=X\setminus\bigcap_{a\in\mathcal{A}}\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]$
$=\bigcup_{a\in\mathcal{A}}X\setminus\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]=\bigcup_{a\in\mathcal{A}}a$
The last inclusion holds as for each $a$, $X\setminus a=\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]$ as $\pi$ is one-to-one. Then $a=X\setminus(X\setminus a)=X\setminus\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]$. It follows that $\bigcup_{a\in\mathcal{A}}a\notin I$. Since $\mathcal{A}\subseteq\mathcal{B}$, we conclude that
$\pi^{-1}[\bigcup_{a\in\mathcal{A}}Y\setminus(\pi^{\prime\prime}[X\setminus a])]=\bigcup_{a\in\mathcal{A}}a\in I^{*}.$
Namely, $\bigcup_{a\in\mathcal{A}}Y\setminus(\pi^{\prime\prime}[X\setminus a])\in\pi_{*}(I^{*})=\pi_{*}(I)^{*}$.
For $(2)$, again, let $\mathcal{B}\subseteq I$ be a cofinal set witnessing that $I$ is deterministic. Consider $\mathcal{C}=\{b\cap A\mid b\in\mathcal{B}\}$. Then $\mathcal{C}$ is cofinal in $I\cap P(A)$. If $\bigcup_{b\in\mathcal{A}}b\cap A\notin I\cap P(A)$, then $\bigcup_{b\in\mathcal{A}}b\cap A\notin I$ (as it is clearly in $P(A)$). It follows that $\bigcup_{b\in\mathcal{A}}b\notin I$ and since $I$ is deterministic, $\bigcup_{b\in\mathcal{A}}b\in I^{*}$.
It follows that $\bigcap_{b\in\mathcal{A}}X\setminus b\in I$ and $A\cap\bigcap_{b\in\mathcal{A}}X\setminus b\in I\cap P(A)$. But
$A\cap\bigcap_{b\in\mathcal{A}}X\setminus b=\bigcap_{b\in\mathcal{A}}A\setminus(b\cap A)=A\setminus\bigcup_{b\in\mathcal{A}}b\cap A.$
Hence $\bigcup_{b\in\mathcal{A}}b\cap A\in(I\cap P(A))^{*}$. ∎
Note that $(2)$ above can be vacuous if $A\in I$, since in that case $I\cap P(A)$ is not proper. So we should at least assume that $A\in I^{+}$. Generally speaking, it is unclear whether an ideal relative to a positive set has the same Tukey-type. However, if the ideal is deterministic, this type does not change:
Fact 2.26.
Suppose that $I$ is an ideal over $X$ and $A\in I^{+}$, then $I\equiv_{T}I\cap P(A)$.
Proof.
Let $\mathcal{B}$ be a witnessing cofinal set for $I$, and let $f:\mathcal{B}\rightarrow I\cap P(A)$ be the map $f(b)=b\cap A$. Then clearly, the map is monotone and its image is the cofinal set $\mathcal{C}$ from the proof of $(2)$ from the previous proposition (and therefore cofinal). To see it is unbounded, suppose that $\bigcup\mathcal{A}\notin I$, then $\bigcup\mathcal{A}\in I^{*}$ and the computation from the previous proposition applies to show that $\bigcup f^{\prime\prime}\mathcal{A}\in(I\cap P(A))^{*}$ and in particular not in $I$. Hence $f$ is unbounded. ∎
Corollary 2.27.
Suppose that $I^{\omega}\equiv_{T}I$, $I$ is deterministic. Then for every ultrafilter $U$ such that $I\subseteq U^{*}$ and $U$ satisfies the $I$-p.i.p, $U\cdot U\equiv_{T}U$.
Proof.
Since $I$ is deterministic, $I^{\omega}\equiv_{T}I\leq_{T}U$. Since $U$ has the $I$-p.i.p we can apply Theorem 2.6 to conclude that $U\cdot U\equiv_{T}U$. ∎
Given any $\{X_{i}\mid i\in N\}$, where each $X_{i}\subseteq P(\omega)$, there is the smallest ideal (might not be proper) that contains all the $X_{i}$. We denote this ideal by $I(\{X_{i}\mid i\in N\})$. It is generated by the sets $\{\bigcup_{i\in M}b_{i}\mid b_{i}\in X_{i},\ M\in[N]^{<\omega}\}$. We can replace each $X_{i}$ be some cofinal set in $B_{i}$ and obtain the same generated ideal.
Theorem 2.28.
Suppose that $I$ is an ideal $I=I({\langle}I_{n}\mid n<\omega{\rangle})$, where $I_{n}\subseteq I$ are deterministic ideals. Then $I$ is deterministic.
Proof.
Let $\mathcal{B}_{n}\subseteq I_{n}$ be a cofinal set witnessing $I_{n}$ being deterministic. Let $\mathcal{B}^{\prime}_{n}=\{b\cup n\mid b\in\mathcal{B}_{n}\}$. Then $I=I({\langle}\mathcal{B}_{n}\mid n<\omega{\rangle})$. Consider the cofinal set $\mathcal{B}$ of all sets of the form $X_{T,(b^{\prime}_{i})_{i\in T}}:=\bigcup_{n\in T}b^{\prime}_{n}$ where $b^{\prime}_{n}\in\mathcal{B}^{\prime}_{n}$ and $T\in[\omega]^{<\omega}$. Then $\mathcal{B}$ is a cofinal set in $I$. Suppose $\bigcup_{j\in S}X_{T_{j},(b^{j}_{i})_{i\in T_{j}}}\notin I$. Note that $X_{T,(b^{\prime}_{i})_{i\in T}}\supseteq\max(T)$. And so, if $\bigcup_{j\in S}T_{j}$ is unbounded in $\omega$, then $\bigcup_{j\in S}X_{T_{j},(b^{j}_{i})_{i\in T_{j}}}=\omega\in I^{*}$. Otherwise, $\bigcup_{j\in S}T_{j}$ is bounded by some $N$ and therefore
$\bigcup_{j\in S}X_{T_{j},(b^{j}_{i})_{i\in T_{j}}}\subseteq\{0,...,N\}\cup\bigcup_{i\leq N}(\bigcup_{b\in\mathcal{A}_{i}}b),$
where $\mathcal{A}_{i}=\{b^{i}_{j}\mid i\in T_{j}\}\subseteq\mathcal{B}_{i}$. Since this is a finite union of sets which is not in $I$, there is $i\leq N$ such that $\bigcup_{b\in\mathcal{A}_{i}}b\notin I$, and in particular, $\bigcup_{b\in\mathcal{A}_{i}}b\notin I_{i}$. Since $\mathcal{B}_{i}$ is a witness for $I_{i}$ being deterministic, $\bigcup_{b\in\mathcal{A}_{i}}b\in I_{i}^{*}\subseteq I^{*}$. Since $\bigcup_{b\in\mathcal{A}_{i}}b\subseteq\bigcup_{j\in S}X_{T_{j},(b^{j}_{i})_{i\in T_{j}}}$, it follows that $\bigcup_{j\in S}X_{T_{j},(b^{j}_{i})_{i\in T_{j}}}\in I^{*}$. ∎
Proposition 2.29.
Suppose that $\mathrm{fin}\subseteq I$ be any ideal over $\omega$, and ${\langle}J_{n}\mid n<\omega{\rangle}$ is a sequence of deterministic ideals over $\omega$ such that for every $n<\omega$, $J_{n+1}\geq_{T}J_{n}$. Then $\sum_{I}J_{n}$ is deterministic.
Proof.
Let $\mathcal{B}_{n}\subseteq J_{n}$ witness that $J_{n}$ is deterministic. Let ${\langle}f_{m,n}:J_{n}\rightarrow J_{m}\mid n\leq m<\omega{\rangle}$ be a sequence of unbounded maps. Denote by $\mathcal{B}$ the set of all sequences $\vec{b}={\langle}b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n<m<\omega$, $b_{m}\supseteq f_{m,n}(b_{n})$. For $\vec{b}\in\mathcal{B}$ and $A\in I$ defined
$C_{A,\vec{b}}=\bigcup_{n\in A}\{n\}\times\omega\cup\bigcup_{n\notin A}\{n\}\times b_{n}.$
Let us show that $\mathcal{C}=\{C_{A,\vec{b}}\mid A\in I,\vec{b}\in\mathcal{B}\}$ is cofinal in $\sum_{I}J_{n}$. Let $Z\in\sum_{I}J_{n}$, and $A=\{n\mid(Z)_{n}\notin J_{n}\}$. By definition of $\sum_{I}J_{n}$, $A\in I$. Construct an increasing sequence $\langle b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n\notin A$, $(Z)_{n}\subseteq b_{n}$ and for every $n<m<\omega$, $f_{m,n}(b_{n})\subseteq b_{m}$. It is possible to construct such a sequence recursively. At stage $n$, we need to find $b_{n}$ such that for all $k<n$, $f_{n,k}(b_{k})\subseteq b_{n}$ and if $n\notin A$, then also $(Z)_{n}\subseteq b_{n}$. These are finitely many sets in $J_{n}$ and therefore, since $\mathcal{B}_{n}$ is a cofinal set in $J_{n}$, we can find a single $b_{n}\in\mathcal{B}_{n}$ including the union of these sets. It follows that for every $n$, $(Z)_{n}\subseteq(C_{A,{\vec{b}}})_{n}$ and therefore $Z\subseteq C_{A,\vec{b}}$. Let us prove that $\mathcal{C}$ witnesses that $\sum_{I}J_{n}$ is deterministic. Suppose that $\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}}\notin\sum_{I}J_{n}$. Then $A:=\{n<\omega\mid(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}\notin J_{n}\}\notin I$. Note that $(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}$. If there is $i_{0}\in T$ such that $n\in X_{i_{0}}$, then $(C_{X_{i_{0}},\vec{b}^{i_{0}}})_{n}=\omega$ and in particular $\omega=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}$. Otherwise, consider $n\in A$ such that for every $i\in T$, $n\notin X_{i}$. Then $(C_{X_{i},\vec{b}^{i}})_{n}\in\mathcal{B}_{n}$ for every $i\in T$ and $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\notin J_{n}$. Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J^{*}_{n}$. Fix any $n_{0}\in A$, which exists  [^17] If $I$ is principle then we pick $n_{0}$ such that $\{n_{0}\}\in I^{+}$. as $A\notin I$. For every $n\geq n_{0}$, either there is $i\in T$ such that $n\in X_{i}$, and as we have seen, $\cup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}=\omega\in J^{*}_{n}$. Otherwise, for every $i\in T$, $(C_{X_{i},\vec{b}^{i}})_{n}=b^{i}_{n}\in\mathcal{B}_{n}$, and by the assumption, $f_{n,n_{0}}(b^{i}_{n_{0}})\subseteq b^{i}_{n}$. Since $n_{0}\in A$, $\bigcup_{i\in T}b^{i}_{n_{0}}\notin J_{n_{0}}$, and since $f_{n,n_{0}}$ is unbounded, $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\notin J_{n}$. Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, it follows that $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\in J^{*}_{n}$. We conclude that for every $n\geq n_{0}$, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J^{*}_{n}$. Since $\{n_{0},n_{0}+1,...\}\in I^{*}$, we have that $\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}}\in(\sum_{I}J_{n})^{*}$ as wanted.
Corollary 2.30.
Suppose that $I,J$ are ideals over $\omega$ such that $J$ is deterministic. Then $I\otimes J$ is deterministic.
Corollary 2.31.
For every $I$, $I\otimes\mathrm{fin}$ is deterministic.
Corollary 2.32.
For any non-principal ultrafilter $U$ such that $U\equiv_{T}U\cdot U$ and for every $I\subseteq U^{*}$, $I^{\omega}\leq_{T}U$.
Proof.
$I\leq_{T}I\otimes\mathrm{fin}\leq_{T}U\cdot U$. Hence by Corollary 1.8
$I^{\omega}\leq_{T}(U\cdot U)^{\omega}\equiv_{T}(U\cdot U)\cdot(U\cdot U)\equiv_{T}U\cdot U\equiv_{T}U.$
3. Proof of the main result and some corollaries
Let us turn to the proof of the main result which appears in Theorem 3.2 below.
Corollary 3.1.
For any non-principal ultrafilter $U,V$ over $\omega$, $V\cdot U,\ U\cdot V\geq_{T}(U^{*}\cap V^{*})^{\omega}$
Proof.
We have that $U^{*}\cap V^{*}\subseteq V^{*}$ and $\mathrm{fin}\subseteq V^{*}$. We conclude that $(U^{*}\cap V^{*})\otimes\mathrm{fin}\subseteq(V\cdot V)^{*}$. By Corollary 2.31, $(U^{*}\cap V^{*})\otimes\mathrm{fin}$ is deterministic, hence
$U^{*}\cap V^{*}\leq_{T}(U^{*}\cap V^{*})\otimes\mathrm{fin}\leq_{T}V\cdot V$
By Milovich’s Corollary 1.8 and Theorem 1.7
$U\cdot V\equiv_{T}U\cdot(V\cdot V)\equiv_{T}U\times(V\cdot V)^{\omega}\geq_{T}(U^{*}\cap V^{*})^{\omega}.$
The proof that $V\cdot U\geq_{T}(U^{*}\cap V^{*})^{\omega}$ is symmetric. ∎
Theorem 3.2.
Suppose that $U,V$ are ultrafilters on $\omega$, then
$U\cdot V\equiv_{T}U\times V\times(U\cap V)^{\omega}.$
Hence $U\cdot V\equiv_{T}V\cdot U$.
Proof.
By the previous corollary, $(U\cap V)^{\omega}\equiv_{T}(U^{*}\cap V^{*})^{\omega}\leq_{T}U\cdot V$. Since $U\times V\times(U\cap V)^{\omega}$ is the least upper bound in the Tukey order of $U,V,(U\cap V)^{\omega}$, we have
$U\times V\times(U\cap V)^{\omega}\leq_{T}U\cdot V$
For the other direction, recall that by Theorem 2.13 $V$ has the $(U\cap V)^{*}$-p.i.p and therefore by Proposition 2.5
$U\cdot V\equiv_{T}U\times V^{\omega}\leq_{T}U\times V\times(U^{*}\cap V^{*})^{\omega}\equiv_{T}U\times V\times(U\cap V)^{\omega}.$
Recall that for an ultrafilter $U$, $Ch(U)$ denoted the minimal size of a cofinal subset of $U$. It is clear that if $U\equiv_{T}U^{\prime}$ then $Ch(U)=Ch(U^{\prime})$. As a corollary of theorem 3.2, we get:
Corollary 3.3.
For any two ultrafilters, $Ch(U\cdot V)=Ch(V\cdot U)$.
It is possible to prove by induction now that the product of $n$-ultrafilters commute, but we would like to get the exact cofinal type of such product. We will need the following facts:
Fact 3.4.
  1. (1)
    $U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times(U_{2})^{\omega}...\times(U_{n})^{\omega}$.
  2. (2)
    $(\mathbb{P}^{\omega})^{\omega}\simeq\mathbb{P}^{\omega}$.
Proof.
$(1)$ is just by induction using Milovich Theorem 1.7. For $(2)$, just decompose $\omega$ into infinitely many infinite sets. See for example [3, Fact 2.4]. ∎
Corollary 3.5.
For every ultrafilters $U_{1},...,U_{n}$,
$U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times U_{2}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
Proof.
Following the proof of Theorem 3.2, $U_{1}\cap...\cap U_{n}\subseteq U_{n}$ and we can prove similarly that $(U_{1}\cap....\cap U_{n})^{\omega}\leq_{T}U_{n}\cdot U_{n}\leq_{T}U_{1}\cdot...\cdot U_{n}$. We conclude that
$U_{1}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}\leq_{T}U_{1}\cdot...\cdot U_{n}$
In the other direction, each $U_{i}$ has the $(U_{1}\cap...\cap U_{n})^{*}$-p.i.p and therefore
$U_{1}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times(U_{2})^{\omega}\times(U_{3})^{\omega}\times...\times(U_{n})^{\omega}\leq_{T}$
$\leq_{T}U_{1}\times U_{2}\times(U_{1}\cap....\cap U_{n})^{\omega}\times U_{3}\times(U_{1}\cap...\cap U_{n})^{\omega}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
$\leq_{T}U_{1}\times...\times U_{n}\times((U_{1}\cap...\cap U_{n})^{\omega})^{\omega}\equiv_{T}U_{1}\times....\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
Corollary 3.6.
Suppose that $U\cdot U\equiv_{T}U$ then for every $V\geq_{T}U$, $V\cdot V\equiv_{T}V$. Namely the class of ultrafilters $U$ such that $U\cdot U\equiv_{T}U$ is upward closed and the class of ultrafilters $V$ such that $V\cdot V>_{T}V$ is downward closed.
Proof.
Otherwise, $V<_{T}V\cdot V$ and
$U\cdot V\equiv_{T}U\times V^{\omega}\equiv_{T}V^{\omega}\equiv_{T}V\cdot V>_{T}V\equiv_{T}V\times U\equiv_{T}V\times U^{\omega}\equiv_{T}V\cdot U,$
contradiction. ∎
Corollary 3.7.
For every $\alpha$, $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
Proof.
The base case $\mathrm{fin}^{\otimes 1}=\mathrm{fin}$ which is deterministic. Suppose this is true for $\alpha$, then $\mathrm{fin}^{\otimes\alpha+1}=\mathrm{fin}\cdot\mathrm{fin}^{\otimes\alpha}$ and by induction hypothesis and Corollary 2.30 $\mathrm{fin}^{\otimes\alpha+1}$ is deterministic. At limit step, $\mathrm{fin}^{\otimes\alpha}=\sum_{\mathrm{fin}}\mathrm{fin}^{\otimes\alpha_{n}}$. By it is known that $\mathrm{fin}^{\alpha_{n}}$ is increasing in the Tukey order (even in the Rudin-Keisler order, see for example [3, Lemma 3.2]) and by induction hypothesis are all deterministic. Therefore, by Proposition 2.29 $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
In [3] it was noticed that a generic ultrafilter for $P(X)/I$ where $I$ is a $\sigma$-ideal satisfies the $I$-p.i.p. now together with Corollary 2.27, we recover the result from [3] in our abstract settings:
Corollary 3.8.
For every $\alpha$, a generic ultrafilter $G$ for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfies $G\cdot G\equiv_{T}G$.
Proof.
$G$ has the $\mathrm{fin}^{\otimes\alpha}$-p.i.p, and by Corollary 3.7 $\mathrm{fin}^{\otimes\alpha}$ is deterministic. Also note that $(\mathrm{fin}^{\otimes\alpha})^{\omega}\equiv_{T}\mathrm{fin}^{\otimes\alpha}$ (see [3, Thm. 3.3 & Fact 2.4]). Thus by Corollary 2.27 $G\cdot G\equiv_{T}G$. ∎
4. On ultrafilters above $I^{\omega}$
The cofinal type of $\omega^{\omega}$ came up in several early papers [28, 35] regarding the Tukey order on ultrafilters. It turns out that some general problems boil down to being Tukey above $\omega^{\omega}$. Such a problem is addressed in Theorem 0.2, which sets up the equivalence for $p$-point ultrafilter $U$, between $U\cdot U\equiv_{T}U$ and $U\geq_{T}\omega^{\omega}$. This was genralized in Theorem 0.3 which ensures that for a general ultrafilter $U$, $U\cdot U\equiv_{T}U$ is equivalent to the existence of some $I\subseteq U^{*}$ such that $U$ has the $I$-p.i.p and $U\geq_{T}I^{\omega}$. In the first part of this section, we tighten the connection between the $I$-p.i.p and being above $I^{\omega}$ for a deterministic ideal $I$. Then, in the second subsection, we shall restrict our attention to $I=\omega$.
4.1. The case of a deterministic ideal $I$
There is a slight difference between the type of equivalence in Theorem 0.2 to the equivalence $U\cdot U\equiv_{T}U$ for $p$-points and the general one in 0.3. Indeed, in the latter, the ideal $I$ can vary. Hence it is unclear in general if for a fixed $I$, the following is true: for any ultrafilter $U$ which has the $I$-p.i.p, $U\cdot U\equiv_{T}U$ iff $U\geq_{T}I^{\omega}$. Let us note first that such equivalence holds for simple ideals (and therefore also for deterministic ideals).
Proposition 4.1.
Suppose that $I$ is simple. Then for any ultrafilter $U$ which has the $I$-p.i.p, $U\cdot U\equiv_{T}U\times I^{\omega}$. Therefore, the following are equivalent:
  1. (1)
    $U\cdot U\equiv_{T}U$.
  2. (2)
    $U\geq_{T}I^{\omega}$.
Proof.
By Theorem 1.7, $U\cdot U\equiv_{T}U^{\omega}$. Since $I$ is simple, and $I\subseteq U^{*}$, $U\geq_{T}I$ and in particular $U\cdot U\equiv_{T}U^{\omega}\geq_{T}U\times I^{\omega}$. The other direction follows from the $I$-p.i.p of $U$ and Proposition 2.5. Now to see the equivalence, $(2)\Rightarrow(1)$ follows from Theorem 2.6, and $(1)\Rightarrow(2)$ follows from the first part as $U\cdot U\equiv_{T}U\times I^{\omega}\leq_{T}U\leq_{T}U\cdot U$. ∎
Our next objective it to study the class of ultrafilters which are Tukey above $I^{\omega}$. The next theorem shows that for deterministic $I$’s this class extends the class of ultrafilters which do not have the $I$-p.i.p.
Theorem 4.2.
Suppose that $\mathrm{fin}\subseteq I\subseteq U^{*}$ is deterministic, then if $U\not\geq_{T}I^{\omega}$, then $U$ has the $I$-p.i.p
Proof.
Let us verify the equivalent condition in Proposition 2.7, let ${\langle}A_{n}\mid n<\omega{\rangle}$ be a partition of $\omega$ such that $A_{n}\notin U$. We need to find $X\in U$ such that $X\cap A_{n}\in I$ for every $n$. Without loss of generality, suppose that $A_{n}\in I^{+}$ for every $n$. Since $\mathrm{fin}\subseteq I$, $A_{n}$ is infinite and we can find a bijection $\pi:\omega\leftrightarrow\omega\times\omega$ such that $\pi^{\prime\prime}A_{i}=\{i\}\times\omega$. Let $W=\pi_{*}(U)$ be the Rudin-Keisler isomorphic copy of $U$. For each $n<\omega$, consider the ideal $I_{n}=\pi_{*}(I\cap P(A_{n}))$ on $\{n\}\times\omega$. By Proposition 2.25, $I\cap P(A_{n})$ is a deterministic and since $\pi\restriction A_{n}$ is one-to-one $I_{n}=\pi_{*}(I\cap P(A_{n}))$ is deterministic. It follows by 2.29 $\sum_{\mathrm{fin}}I_{n}$ is deterministic. Moreover, by Fact 2.26, $I\equiv_{T}I\cap P(A_{n})\equiv_{T}I_{n}$ and therefore
$I^{\omega}\equiv_{T}\prod_{n<\omega}I_{n}\equiv_{T}\sum_{\mathrm{fin}}I_{n}$
Since $U\not\geq_{T}I^{\omega}$, $W\not\geq_{T}\sum_{\mathrm{fin}}I_{n}$. Since $\sum_{\mathrm{fin}}I_{n}$ is deterministic, it follows that $\sum_{\mathrm{fin}}I_{n}\not\subseteq W^{*}$. Thus, there is $X^{\prime}\in\sum_{\mathrm{fin}}I_{n}\cap W$. Namely, for all but finitely many $n$’s, $(X^{\prime})_{n}\in I_{n}$. Since each $A_{i}\notin U$, we may assume that for every $n$, $(X^{\prime})_{n}\in I_{n}$. Let $X=\pi^{-1}[X^{\prime}]$, then for every $n<\omega$, $X\cap A_{n}\in I$ as $\pi^{\prime\prime}X\cap A_{n}=\{x\}\cap(X)_{n}\in I_{n}$. ∎
The proof of the above actually gives the following:
Corollary 4.3.
Suppose that $U$ does not have the $I$-p.i.p, then there is $W\subseteq\omega\times\omega$ such that $W\equiv_{RK}U$ and $\sum_{\mathrm{fin}}I_{n}\subseteq W$, and each $I_{n}\equiv_{T}I$. In particular, the Tukey type of $I^{\omega}$ is realized as a sub-ideal of $U^{*}$.
Taking $I=\mathrm{fin}$ in the above we obtain the following corollary
Corollary 4.4.
Suppose that $U$ is a non-principal ultrafilter such that $U\not\geq_{T}\omega^{\omega}$ then $U$ is a $p$-point.
As a corollary, we see that in Proposition 4.1 and therefore also in Theorem 0.2, the $I$-p.i.p assumption is somewhat redundant.
Corollary 4.5.
If $I\subseteq U^{*}$ is deterministic then the following are equivalent:
  1. (1)
    $U\equiv_{T}U\cdot U$ or $U$ does not have the $I$-p.i.p.
  2. (2)
    $U\geq_{T}I^{\omega}$
4.2. Ultrafilters above $\omega^{\omega}$
As observed by Dobrinen and Todorcevic, rapid ultrafilters form a subclass of those which are Tukey above $\omega^{\omega}$. By Miller [26], in the Laver model where the Borel conjecture holds181818 i.e. the model obtained by adding $\omega_{2}$-many Laver reals to a model of $CH$., there are no rapid ultrafilters. On the other hand, there are always ultrafilters which are above $\omega^{\omega}$, namely tukey-top ultrafilters. The question regarding the existence (in $ZFC$) of non-Tukey top ultrafilters is a major open problem. Hence we do not expect to have a simple $ZFC$ construction of a non-Tukey-top ultrafilter which is above $\omega^{\omega}$.
The result from the previous section entails some drastic consistency results regarding the class of ultrafilters above $\omega^{\omega}$:
Corollary 4.6.
Suppose that there are no $p$-points, then every ultrafilter is above $\omega^{\omega}$.
Models with no $p$-points were first constructed by Shelah [34] and later by Chudonsky and Guzman [11].
By yet another result of Shelah, in the Miller model [27], which is obtained by countable support iteration of the superperfect tree forcing of length $\omega_{2}$ over a model of CH, every $p$-point is generated by $\aleph_{1}$-many sets. It is known that $\mathfrak{d}=\mathfrak{c}$ holds in that model. Therefore, every $p$-point is generated by less than $\mathfrak{d}$-many sets and in particular not above $\omega^{\omega}$.
Corollary 4.7.
It is consistent that $p$-points are characterized by not being above $\omega^{\omega}$.
Focusing on models less extreme than the ones above, we may be interested in those $p$-point which are above $\omega^{\omega}$. The purpose of this section is to address the question raised in [3] whether rapid $p$-point are exactly those $p$-points which are above $\omega^{\omega}$ (The dashed area in Figure 4.2). As a warm-up, let us note that there are always non-rapid ultrafilters above $\omega^{\omega}$. To see this, we need the following result [26, Thm. 4]:
Proposition 4.8 (Miller).
For any two ultrafilters $U,V$ on $\omega$, $U\cdot V$ is rapid iff $V$ is rapid.
By results of Choquet [12], there are always non-rapid ultrafilters. Taking any such $U$, $U\cdot U$ is certainly above $\omega^{\omega}$ and by the above result of Miller, it is non-rapid.
Corollary 4.9.
There is a non-rapid ultrafilter which is Tukey above $\omega^{\omega}$.
Note that the ultrafilter we constructed in the previous proof in not a $p$-point as it is a product.
[…]
The real issue is to construct a $p$-point which is not rapid but still above $\omega^{\omega}$. To do that, let us introduce the class of $\alpha$-almost rapid ultrafilters.
Given a function $f:\omega\rightarrow\omega\setminus\{0\}$ such that $f(0)>0$. We denote by $exp(f)(0)=f(0)$ and
$exp(f)\big{(}n+1\big{)}=f\big{(}exp(f)(n)\big{)}=f(f(f(f...f(0)..))).$
We define the $n^{\text{th}}$ $f$-exponent function,
$exp_{0}(f)=f\text{ and }exp_{n}(f)=exp(exp_{n-1}(f)).$
This definition continuous transfinitely for every $\alpha<\omega_{1}$:
$exp_{\alpha+1}(f)=exp(exp_{\alpha}(f)).$
For limit $\delta<\omega_{1}$, we fix some increasing cofinal sequence ${\langle}\delta_{n}\mid n<\omega{\rangle}$ in $\delta$, and let
$exp_{\delta}(f)(n)=\max\{exp_{\delta_{n}}(f)(n),exp_{\delta}(f)(n-1)+1\}.$
Lemma 4.10.
Let $f,g:\omega\rightarrow\omega$ be increasing functions.
  1. (1)
    For every $\alpha<\omega_{1}$, $exp_{\alpha}(f)$ is increasing.
  2. (2)
    If $f\leq g$ then for every $\alpha<\omega_{1}$, $exp_{\alpha}(f)\leq exp_{\alpha}(g)$.
  3. (3)
    For every $\alpha<\beta<\omega_{1}$, $exp_{\alpha}(f)<^{*}exp_{\beta}(f)$.
Proof.
For $(1)$, we proceed by induction. For $\alpha=0$, $exp_{0}(f)=f$ is increasing. Suppose $exp_{\alpha}(f)$ is increasing, then for every $n<\omega$, $exp_{\alpha}(f)(n)>n$. For $\alpha+1$, let $n<\omega$. Since $exp_{\alpha}(f)$ is increasing,
$exp_{\alpha+1}(f)(n+1)=exp_{\alpha}(f)(exp_{\alpha+1}(f)(n))>exp_{\alpha+1}(f)(n).$
For limit $\delta$, is clear from the definition that $exp_{\delta}(f)$ is increasing. Also (2) is proven by induction. The base case is $exp_{0}(f)=f\leq g=exp_{0}(g)$. Suppose this was true for $\alpha$, and let us prove by induction on $n<\omega$. The base again is
$exp_{\alpha+1}(f)(0)=exp_{\alpha}(f)(0)\leq exp_{\alpha}(g)(0)\leq exp_{\alpha+1}(g)(0)$
Suppose that $exp_{\alpha+1}(f)(n)\leq exp_{\alpha+1}(g)(n)$, then by $(1)$ and the induction hypothesis
$exp_{\alpha+1}(f)(n+1)=exp_{\alpha}(f)(exp_{\alpha+1}(f)(n))\leq exp_{\alpha}(f)(exp_{\alpha+1}(g)(n))$
$\leq exp_{\alpha}(g)(exp_{\alpha+1}(g)(n))=exp_{\alpha+1}(g)(n+1)$
At limit stages $\delta$, by the induction hypothesis,
$exp_{\delta}(f)(n)=exp_{\delta_{n}}(f)(n)\leq exp_{\delta_{n}}(g)(n)=exp_{\delta}(g)(n).$
Finally, $(3)$ is a standard diagonalization argument. ∎
Definition 4.11.
For $\alpha<\omega_{1}$, we say that an ultrafilter $U$ is $\alpha$-almost-rapid if for every function $f\in\omega^{\omega}$ there is $X\in U$ such that $exp_{\alpha}(f_{X})\geq^{*}f$, where $f_{X}$ is the increasing enumeration of $X$.
Remark 4.12.
By strengthening the above definition, we may require that $exp_{\alpha}(f_{X})\geq f$. However, this strengthening turns out to be an equivalent definition.
Note that 00-almost-rapid is just rapid, and by $(3)$ of the previous lemma, if $\beta\leq\alpha$ then $\beta$-almost-rapid implies $\alpha$-almost-rapid. We call $U$ almost-rapid if it is $1$-almost-rapid.
Proposition 4.13.
If $U$ is $\alpha$-almost-rapid implies $U\geq_{T}\omega^{\omega}$
Proof.
Consider the map $X\mapsto exp_{\alpha}(f_{X})$. We claim that it is monotone and cofinal. First, suppose that $X\subseteq Y$, then the natural enumerations $f_{X},f_{Y}$ of $X,Y$ (resp.) satisfy $f_{X}\geq f_{Y}$. Then by Lemma 4.10(3) $exp_{\alpha}(f_{X})\geq exp_{\alpha}(f_{Y})$. The map above is cofinal by the $\alpha$-almost rapidness of $U$. ∎
Rapid ultrafilters are characterized by the following property [26]: An ultrafilter over $\omega$ is rapid if and only if for every sequence ${\langle}P_{n}\mid n<\omega{\rangle}$ of finite subsets of $\omega$, there is $X\in U$ such that for every $n<\omega$, $|X\cap P_{n}|\leq n$. The proposition below provides an analogous characterization of almost-rapid ultrafilters.
Proposition 4.14.
The following are equivalent:
  1. (1)
    $U$ is almost-rapid.
  2. (2)
    For any sequence ${\langle}P_{n}\mid n<\omega{\rangle}$ of sets, such that $P_{n}$ is finite, there is $X\in U$ such that for each $n<\omega$, $exp(f_{X})(n-1)\geq|X\cap P_{n}|$ (where $exp(f_{X})(-1)=0$).
Proof.
$(1)\Rightarrow(2)$: Suppose that $U$ is almost rapid, and let ${\langle}P_{n}\mid n<\omega{\rangle}$ be a sequence as above. Let $f(n)=\max(P_{n})+1$. By $(1)$, there is $X\in U$ which is obtained by almost-rapidness, namely $exp(f_{X})(0)=f_{X}(0)=\min(X)>f(0)>\max(P_{0})$ and therefore $X\cap P_{0}=\emptyset$. Next, $exp(f_{X})(1)=f_{X}(f_{X}(0))>f(1)>\max(P_{1})$ hence $|P_{1}\cap X|\leq f_{X}(0)=exp(f_{X})(0)$. In general $f_{X}(exp(f_{X})(n))=exp(f_{X})(n+1)>f(n+1)>\max(P_{n+1})$ and therefore $|X\cap P_{n+1}|\leq exp(f_{X})(n)$.
$(2)\Rightarrow(1)$: Let $f$ be any function. Let $P_{n}=f(n)$. Then by $(2)$, there is $X$ such that $|X\cap f(n)|\leq exp(f_{X})(n-1)$. In particular, $X\cap f(0)=\emptyset$ and therefore $exp(f_{X})(0)=f_{X}(0)=\min(X)\geq f(0)$. In general, $|X\cap f(n)|\leq exp(f_{X})(n-1)$ and therefore $f(n)<f_{X}(exp(f_{X})(n-1))=exp(f_{X})(n)$. ∎
Theorem 4.15.
Assume CH. Then there is a $p$-point which is almost-rapid but not rapid
Proof.
Let $P_{n}=\{1,...,2^{n}\}$. Let $I=\{A\subseteq\omega\mid\exists k\forall n,\ |A\cap P_{n}|\leq k\cdot n\}$. Then $I$ is a proper ideal on $\omega$. Suppose that ${\langle}P_{n}\mid n<\omega{\rangle}$ is not a counterexample for $U$ being rapid, then there is a set $X\in U$ such that $|X\cap P_{n}|\leq n$ for every $n$ and therefore $X\in I$. Hence, as long as we have $U\subseteq I^{+}$, $U$ will not be rapid. Note that
$A\in I^{+}\text{ iff for every }k,\ \text{there is }n_{k}\text{ such that }|A\cap P_{n_{k}}|>kn_{k}.$
Or equivalently, $n\mapsto|A\cap P_{n}|$ is not asymptotically bounded by a linear function of $n$. The following is the key lemma for our construction:
Lemma 4.16.
Suppose that ${\langle}A_{n}\mid n<\omega{\rangle}\subseteq I^{+}$ is $\subseteq$-decreasing, and $f:\omega\rightarrow\omega$. Then there is $B\subseteq\omega$ such that
  1. (1)
    $B\subseteq^{*}A_{n}$ for every $n$.
  2. (2)
    $B\in I^{+}$.
  3. (3)
    $exp(f_{B})>f$.
Proof.
Suppose without loss of generality that $f$ is increasing. In particular, $f(k)\geq k$. Consider $f(1)$, find $2<n_{1}$ so that
$|A_{1}\cap P_{n_{1}}|>(f(1)+2)\cdot n_{1}$
such an $n_{1}$ exists as $A_{1}$ is positive and taking $k=f(1)+2$. Find $a_{0},...,a_{n_{1}+1}$ such that
  1. (1)
    $f(0),n_{1}+1<a_{0}<a_{1}<....<a_{n_{1}+1}$.
  2. (2)
    $a_{n_{1}+1}>f(1)$.
  3. (3)
    $a_{0},a_{1},...,a_{n_{1}+1}\in A_{1}\cap P_{n_{1}}$.
It is possible to find such elements as
$|A_{1}\cap P_{n_{1}}\setminus\{0,...,n_{1}+1\}|>(f(1)+2)n_{1}-(n_{1}+2)\geq 3n_{1}-n_{1}-2=2n_{1}-2\geq n_{1}+1.$
So there are $n_{1}+1$ elements in $A_{1}\cap P_{n_{1}}$ greater than $n_{1}+1$. Since $|A_{1}\cap P_{n_{1}}|>f(1)$, we can also make sure that the $n_{1}+1$ element we choose is above $f(1)$. This way, we have guaranteed that:
  1. (1)
    $f(0)<f(1)<a_{0}$.
  2. (2)
    $a_{a_{0}}$ was not defined yet (!), but as long as the sequence is increasing, $a_{a_{0}}>f(1)$.
  3. (3)
    For $k=1$, there is $n_{1}$ such that $|B\cap P_{n_{1}}|>n_{1}$
Now consider $f(2)$ and find $n_{2}>2,a_{n_{1}+1}$ so that
$|A_{2}\cap P_{n_{2}}|>(f(2)+2)(a_{n_{1}+1}+1)n_{2},$
we find
  1. (1)
    $n_{1}+1+2n_{2}+1<a_{n_{1}+2}<...<a_{n_{1}+1+2n_{2}+1}$.
  2. (2)
    $f(2)<a_{n_{1}+1+2n_{2}+1}$.
  3. (3)
    $a_{n_{1}+2},...,a_{n_{1}+1+2n_{2}+1}\in A_{2}\cap P_{2}$.
This is possible to do since
$|A_{2}\cap P_{n_{2}}\setminus\{0,...,n_{1}+2n_{2}+2\}|>(f(2)+2)(a_{n_{1}+1}+2)n_{2}-(n_{1}+2n_{2}+3)>$
$>8n_{2}-(3n_{2}+3)=6n_{2}-3>2n_{2}+1$
So we can find $a_{n_{1}+2},...,a_{n_{1}+1+2n_{2}+1}$ above $n_{1}+1+2n_{2}+1$ (and therefore also above $a_{n_{1}+1}$). We can also make sure that the last element we pick is above $f(2)$. This way we ensured the following:
  1. (1)
    As we observed, $a_{a_{0}}$ was not defined in the first round (and might not be defined in the second round as well) and therefore (a possibly future) $a_{1}^{\prime}:=a_{a_{0}}>n_{1}+1+2n_{2}+1$. Thus a future $a^{\prime}_{2}:=a_{a_{a_{0}}}>a_{n_{1}+1+2n_{2}+1}>f(2)$.
  2. (2)
    For $k=2$, there is $n_{2}$ such that $|B\cap P_{n_{2}}|>2n_{2}$.
In general suppose we have defined $n_{1}<n_{2}<....<n_{k}$ and $a_{0},...,a_{\sum_{i=1}^{k}in_{i}+1}$ and such that $a^{\prime}_{k-1}>\sum_{i=1}^{k}in_{i}+1$. Then we find $n_{k+1}>k+1,a_{\sum_{i=1}^{k}in_{i}+1}$ such that $|A_{k+1}\cap P_{n_{k+1}}|>3(k+1)(f(k+1)+1)n_{k+1}$. We now define
$a_{(\sum_{i=1}^{k}in_{i}+1)+1},a_{(\sum_{i=1}^{k}in_{i}+1)+2}...,a_{\sum_{i=1}^{k+1}in_{i}+1}$
(that is $(k+1)n_{k+1}+1$ many elements) so that:
  1. (1)
    $\sum_{i=1}^{k+1}in_{i}+1<a_{(\sum_{i=1}^{k}in_{i}+1)+1}<...<a_{\sum_{i=1}^{k+1}in_{i}+1}$,
  2. (2)
    $a_{\sum_{i=1}^{k+1}in_{i}+1}>f(k+1)$.
  3. (3)
    $a_{(\sum_{i=1}^{k}in_{i}+1)+1},...,a_{\sum_{i=1}^{k+1}in_{i}+1}\in A_{k+1}\cap P_{n_{k+1}}$.
To see that such $a$’s exists, note that
$|A_{k+1}\cap P_{n_{k+1}}\setminus\{0,...,\sum_{i=1}^{k+1}in_{i}+1\}|>3(k+1)(f(k+1)+1)n_{k+1}-(\sum_{i=1}^{k+1}in_{i}+1)-1$
$>3(k+1)(f(k+1)+1)n_{k+1}-((k+1)n_{k+1}+1)-(\sum_{i=1}^{k}in_{i}+1)-1$
$(k+1)(3f(k+1)+3)n_{k+1}-2((k+1)n_{k+1}+1)$
$>(k+1)(3f(k+1)+1)n_{k+1}>(k+1)n_{k+1}+1$
Hence we can find $(k+1)n_{k+1}+1$-many elements in $A_{k+1}\cap P_{k+1}$ above $\sum_{i=1}^{k+1}in_{i}+1$. Also, since $|A\cap P_{n}|>f(k+1)$ we can make sure that $a_{\sum_{i=1}^{k+1}in_{i}+1}>f(k+1)$. This way we ensure that:
  1. (1)
    Since $a_{a^{\prime}_{k-1}}$ was not previously defined in previous rounds, $a^{\prime}_{k}:=a_{a^{\prime}_{k-1}}>\sum_{i=1}^{k+1}in_{i}+1$ and $a_{a^{\prime}_{k}}$ has not been defined yet. Hence a future $a^{\prime}_{k+1}:=a_{a^{\prime}_{k}}>f(k+1)$.
  2. (2)
    $|B\cap P_{n_{k+1}}|>(k+1)n_{k+1}+1$.
Set $B=\{a_{n}\mid n<\omega\}$. So by the construction, for every $k$ there is $n_{k}$ such that $|B\cap P_{n_{k}}|>kn_{k}$. Hence $B\in I^{+}$. Also, note that $f_{B}(n)=a_{n}$ since the $a_{n}$’s are increasing. By the construction and definition of $exp(f)$, $exp(f_{B})(n)=a^{\prime}_{n}>f(n)$. Finally, note that for each $n$, there is $k$ such that for every $k^{\prime}\geq k$, $a_{k^{\prime}}\in A_{m}$ for some $m\geq n$. Since the sequence of $A_{n}$’s is $\subseteq$-decreasing, $a_{k^{\prime}}\in A_{n}$. We conclude that $B\setminus A_{n}\subseteq\{a_{0},...,a_{k}\}$. ∎
Now for the construction of the ultrafilter. Enumerate $P(\omega)={\langle}X_{\alpha}\mid\alpha<\omega_{1}{\rangle}$, and $P(\omega)^{\omega}={\langle}\vec{A}_{\alpha}\mid\alpha<\omega_{1}{\rangle}$ such that each sequence in $P(\omega)^{\omega}$ appears cofinaly many times in the enumeration. Also enumerate $\omega^{\omega}={\langle}\tau_{\alpha}\mid\alpha<\omega_{1}{\rangle}$. We define a sequence of filters $V_{\alpha}$ such that:
  1. (1)
    $\beta<\alpha\Rightarrow V_{\beta}\subseteq V_{\alpha}$.
  2. (2)
    $V_{\alpha}\subseteq I^{+}$.
  3. (3)
    Either $X_{\alpha}$ or $\omega\setminus X_{\alpha}\in V_{\alpha+1}$.
  4. (4)
    There is $X\in V_{\alpha+1}$ such that $\tau_{\alpha}<exp(f_{X})$.
  5. (5)
    If $\vec{A}_{\alpha}\subseteq V_{\alpha}$ then there is a pseudo-intersection $A\in V_{\alpha+1}$.
Let $V_{0}=I^{*}$. At limit steps $\delta$ we define $V_{\delta}=\bigcup_{\beta<\delta}V_{\beta}$. It is clear that $(1)-(5)$ still holds at limit steps. At successors, given $V_{\alpha}$, since we have only performed countably many steps so far, there are sets $B_{n}\in V_{\alpha}$ such that $V_{\alpha}=I^{*}[{\langle}B_{n}\mid n<\omega{\rangle}]$ where $B_{n}$ is $\supseteq$-decreasing. If either $X_{\alpha}$ or $\omega\setminus X_{\alpha}$ is already in $V_{\alpha}$, we ignore it. Otherwise, we must also have $V_{\alpha}[X_{\alpha}]\subseteq I^{+}$. If $\vec{A}_{\alpha}\not\subseteq V_{\alpha}[X_{\alpha}]$ ignore it. Otherwise, enumerate the set $\{B_{n}\cap X_{\alpha}\mid n<\omega\}\cup\{\vec{A}_{\alpha}(n)\mid n<\omega\}\subseteq V_{\alpha}[X_{\alpha}]$ by ${\langle}B^{\prime}_{n}\mid n<\omega{\rangle}$ and let $C_{n}=\cap_{m\leq n}B^{\prime}_{m}$. We apply the previous lemma to the sequence ${\langle}C_{n}\mid n<\omega{\rangle}$, and $\tau_{\alpha}$ to find $A^{*}\subseteq\omega$ such that:
  1. (1)
    $A^{*}\in I^{+}$.
  2. (2)
    $exp(f_{A^{*}})>\tau_{\alpha}$.
  3. (3)
    $A^{*}\subseteq^{*}C_{n}$ for every $n$.
Since for every $n<\omega$, there is $n^{\prime}$ such that $C_{n^{\prime}}\subseteq\vec{A}_{\alpha}(n)\cap B_{n}$, $A^{*}\subseteq^{*}\vec{A}_{\alpha}(n)$, namely $A^{*}$ is a pseudo intersection of both ${\langle}B_{n}\mid n<\omega{\rangle}$ and $\vec{A}_{\alpha}$. Also, $A^{*}$ is a positive set with respect to the ideal $V_{\alpha}[X_{\alpha}]$. Otherwise, there is some $A\in I^{*}$ and $B_{n}$ such that $A^{*}\cap(A\cap B_{n}\cap X_{\alpha})=\emptyset$. But then $(A^{*}\cap B_{n}\cap X_{\alpha})\cap A=\emptyset$ which implies that $A^{*}\cap B_{n}\cap X_{\alpha}\in I$. However, $A^{*}\subseteq^{*}B_{n}\cap X_{\alpha}$, which implies that $A^{*}\in I$, contradicting property $(1)$ above in the choice of $A^{*}$. Hence we can define $V_{\alpha+1}=V_{\alpha}[X_{\alpha},A^{*}]$ and $(1)-(5)$ hold.
This concludes the recursive definition. The ultrafilter witnessing the theorem is defined by $V^{*}=\cup_{\alpha<\omega_{1}}V_{\alpha}$.
Proposition 4.17.
$V^{*}$ is a non-rapid almost-rapid $p$-point ultrafilter.
Proof.
$V^{*}$ is an ultrafilter since for every $X\subseteq\omega$, there is $\alpha$ such that $X=X_{\alpha}$ and so either $X_{\alpha}$ or $\omega\setminus X_{\alpha}$ are in $V_{\alpha+1}\subseteq V^{*}$. Also $V^{*}$ is a $p$-point since if ${\langle}A_{n}\mid n<\omega{\rangle}\subseteq V^{*}$ then there is $\alpha<\omega_{1}$ such that ${\langle}A_{n}\mid n<\omega{\rangle}\subseteq V_{\alpha}$ and by the properties of the enumeration there is $\beta>\alpha$ such that $\vec{A}_{\beta}={\langle}A_{n}\mid n<\omega{\rangle}$. This means that in $V_{\beta+1}$ there is a pseudo intersection for the $A_{n}$’s. It is non-rapid as $V^{*}\subseteq I^{+}$ and, in fact, the sequence $P_{n}$ witnesses that it is non-rapid. Finally, it is almost rapid since for any function $\tau:\omega\rightarrow\omega$, there is $\alpha$ such that $\tau=\tau_{\alpha}$ and therefore in $V_{\alpha+1}$ there is a set $X$ such that $exp(f_{X})>\tau$. ∎
Corollary 4.18.
It is consistent that there is a $p$-point which is not rapid but still above $\omega^{\omega}$.
Remark 4.19.
CH is not necessary in order to obtain such an ultrafilter, since we can, for example, repeat a similar argument in the iteration of Mathias forcing after we forced the failure of $CH$ and obtain such an ultrafilter. In fact, we conjecture that the construction of Ketonen [22] of a $p$-point from $\mathfrak{d}=\mathfrak{c}$ can be modified to get a non-rapid almost-rapid $p$-point.
5. questions
We collect here some problems which relate to the work of this paper. The first batch of questions regards the Tukey-type of sums of ultrafilters:
Question 5.1.
If $\mathbb{P}$ is below $\mathcal{B}(U,V_{\alpha})$ does it imply that $\mathbb{P}$ is uniformly below $\mathcal{B}(U,V_{\alpha})$? What about the case where $\mathbb{P}$ is an ultrafilter?
Question 5.2.
Is it true in general that $\sum_{U}V_{\alpha}=\inf\mathcal{B}(U,V_{\alpha})$?
Question 5.3.
Is there a nice characterization for the Tukey-type of $\sum_{U}V_{\alpha}$ if we assume that $V_{0}\geq_{T}V_{1}\geq_{T}V_{2}....$?
Question 5.4.
Is there a nice characterization for the Tukey-type of $\sum_{U}V_{\alpha}$ when the sequence of $V_{\alpha}$’s is discrete?
How much of the theory developed here generalizes to measurable cardinals? more concretely:
Question 5.5.
Does Lemma 1.18 hold true for $\sigma$-complete ultrafilters over uncountable cardinals?
The next type of questions relate to the $I$-pseudo intersection property
Question 5.6.
What is the characterization of the $I$-p.i.p property in terms of Skies and Constellations of ultrapowers from [30]?
Question 5.7.
Is the equivalence of Proposition 4.1 true for every ideal $I$?
The following addresses the commutativity of cofinal types between ultrafilters on different cardinals.
Question 5.8.
Is it true that for every two ultrafilters $U,V$ on any cardinals, $U\cdot V\equiv_{T}V\cdot U$?
Let us note that if $U,V$ are $\lambda$ and $\kappa$ ultrafilters respectively, then this holds. To see this, first note that if $\kappa=\lambda\geq_{T}\omega$, then this is a combination of the results from [2] and this paper’s main result. Without loss of generality, $\lambda<\kappa$. In which case, $U\cdot V=U\times V^{\lambda}$. Since $V$ is $\lambda^{+}$-complete, it is not had to see that $V^{\lambda}\equiv_{T}V$, hence $U\cdot V\equiv_{T}V\cdot U$. On the other hand, since $2^{\lambda}<\kappa$, every set $X\in V\cdot U$, contained a set of the form $X\times Y$ for $X\in V$ and $Y\in U$. It follows that also $V\cdot U\equiv_{T}U\times V\equiv_{T}U\cdot V$. So the question above is interesting in case we drop the completeness assumption on the ultrafilters.
Finally, we present a few questions regarding the new class of $\alpha$-almost-rapid ultrafilters.
Question 5.9.
Is it true that for every $\alpha<\beta<\omega_{1}$, the class of $\alpha$-almost-rapid ultrafilters is consistently strictly included in the class of $\beta$-almost-rapid ultrafilters?
We conjecture a positive answer to this question and that similar methods to the one presented in Theorem 4.15 under CH should work.
Question 5.10.
Does $\mathfrak{d}=\mathfrak{c}$ imply that there is a $p$-point which is almost-rapid but not rapid?
Ultimately we would like to understand if such ultrafilters exist in $ZFC$:
Question 5.11.
Is it consistent that there are no almost-rapid ultrafilters?
Following Miller, a natural model would be adding $\aleph_{2}$-many Laver reals. However, the current argument does not immidietly rule out $\alpha$-almost rapid ultrafilters.
Acknowledgment
I would like to thank Gabriel Goldberg for stimulating discussions and for suggesting an example of ultrafilters which eventually led to Proposition 1.21. I would also like to express my gratitude to Natasha Dobrinen for introducing me to the subject, a task which took a great amount of her precious time.

### v2

@@
-The commutativity of cofinal types
+Commutativity of cofinal types
 Abstract.
-We developed the theory of deterministic ideals and present a systematic study of the pseudo-intersection property with respect to an ideal introduced in [3].
-We apply this theory to prove that for any two ultrafilters $U,V$ on $\omega$, $U\cdot V\equiv_{T}V\cdot U$.
-Additionally, we analyze the class of ultrafilters which are Tukey above $\omega^{\omega}$.
-Our theory applies the study of the Tukey types of a general sum of ultrafilters, which, as evidenced by the results of this paper, can be quite complex.
+We continue the study of the pseudo-intersection property with respect to an ideal introduced in [3].
+Our theory applies to the study of the Tukey types of general sums of ultrafilters, which, as evidenced by the results of this paper, can be quite complex.
+It also applies to construct a large class of ultrafilter $\mathcal{C}$ over $\omega$ such that any two ultrafilters $U,V\in\mathcal{C}$ commute; that is, $U\cdot V\equiv_{T}V\cdot U$.
+The class $\mathcal{C}$ class contains most known cofinal types of ultrafilters on $\omega$.
+This is in sharp contrast to the Rudin-Keisler ordering.
+In the third part of this paper, we apply our results to study the class of ultrafilters Tukey above $\omega^{\omega}$.
+Specifically, we prove that ultrafilters without the $I$-p.i.p are always above $I^{\omega}$ and in particular non-$p$-points are Tukey above $\omega^{\omega}$.
 Finally, we introduce the hierarchy of $\alpha$-almost rapid ultrafilters.
-We prove that it is consistent for them to form a strictly wider class than the rapid ultrafilters, and give an example of a non-rapid $p$-point which is Tukey above $\omega^{\omega}$.
-This answers several questions from [2, 3].
-The research of the first author was supported by the National Science Foundation under Grant No. DMS-2246703
+We prove that it is consistent for them to form a strictly wider class than the rapid ultrafilters, and give an example of a non-rapid $p$-point ultrafilter which is Tukey above $\omega^{\omega}$.
+This addresses and answers several questions from [2, 3, 14, 28].
+Key words and phrases:
+ultrafilter, $p$-point, rapid ultrafilter, Tukey order, cofinal type, Fubini product
+2020 Mathematics Subject Classification:
+03E02, 03E04, 03E05, 03E55, 06A06, 06A07
+This research was supported by the National Science Foundation under Grant No. DMS-2346680
 0. Introduction
@@
 Schmidt [33] observed that this is equivalent to having a map $f:P\rightarrow Q$, which is unbounded, namely, $f^{\prime\prime}\mathcal{A}$ is unbounded in $Q$ whenever $\mathcal{A}\subseteq P$ is unbounded in $P$.
-We say that $P$ and $Q$ are Tukey equivalent, and write $P\equiv_{T}Q$, if $P\leq_{T}Q$ and $Q\leq_{T}P$; the equivalence class $[P]_{T}$ is called the Tukey type or cofinal type of $P$.
+We say that $P$ and $Q$ are Tukey equivalent, and write $P\equiv_{T}Q$, if $P\leq_{T}Q$ and $Q\leq_{T}P$; the equivalence class $P]_{T}$ is called the Tukey type or cofinal type of $P$.
 The scope of the study of cofinal types of ultrafilters covers several long-standing open problems such as:
@@
 The Tukey order is also related to the Katovich problem.
-A systematic study of the Tukey order on ultrafilter on $\omega$ traces back to Isbell [19], later to Milovich [28] and Dobrinen and Todorcevic [14].
-Lately, Dobrinen and Benhamou [2] extended this study to ultrafilters on cardinals greater than $\omega$.
-They found a connection to a parallel recent line of research at the level of measurable cardinals, known as the Galvin property studied by Abraham, Benhamou, Garti, Goldberg, Gitik, Hayut, Magidor, Poveda, Shelah and others [1, 15, 16, 6, 4, 5, 8, 7, 17, 9].
-The Galvin property in one of its forms is equivalent to being Tukey top as shown essentially by Isbell (in different terminology).
-Moreover, being Tukey-top in the restricted class of $\sigma$-complete ultrafilters takes the usual studied forms of the Galvin property.
-In this paper, we address the problem of commutativity of the Tukey types of Fubini products of ultrafilters $U,V$ on $\omega$ (Definition 1.1), denoted by $U\cdot V$.
+A systematic study of the Tukey order on ultrafilter over $\omega$, traces back to Isbell [19], later to Milovich [28] and Dobrinen and Todorcevic [14].
+Lately, Benhamou and Dobrinen [2] extended this study to ultrafilters on cardinals greater than $\omega$.
+Over measurable cardinals, the Tukey order is connected to recent developments revolving the so-called Galvin property, studied by Abraham, Benhamou, Garti, Goldberg, Gitik, Hayut, Magidor, Poveda, Shelah and others [1, 15, 16, 6, 4, 5, 8, 7, 17, 9]; the Galvin property in one of its forms is equivalent to being Tukey-top as shown essentially by Isbell (in different terminology).
+Moreover, being Tukey-top in the restricted class of $\kappa$-complete ultrafilters takes the usual studied forms of the Galvin property.
+In this paper, we address the problem of commutativity of the Tukey types of Fubini products of ultrafilters $U,V$ over $\omega$ (Definition 1.1), denoted by $U\cdot V$.
 This problem was suggested in [3], and was already partially addressed:
-  • Dobrinen and Todorchevic [14] proved that if $U,V$ are rapid $p$-points then $U\cdot V\equiv_{T}V\cdot U$.
+  • Dobrinen and Todorcevic [14] proved that if $U,V$ are rapid $p$-points then $U\cdot V\equiv_{T}V\cdot U$.
   • Milovich [29] extended this result to prove that if $U,V$ are just $p$-points, then $U\cdot V\equiv_{T}V\cdot U$.
   • Benhamou and Dobrinen proved later that if $U,V$ are $\kappa$-complete ultrafilters over a measurable cardinal $\kappa$ then $U\cdot V\equiv_{T}V\cdot U$.
-The main result of this paper is:
+Our goal is to study larger classes where commutativity holds, these are called commutative classes of ultrafilters. In that avenue, we prove the following:
 Theorem.
-For any ultrafilters $U_{1},U_{2},...,U_{n}$ on $\omega$, and any permutation $\sigma:\{1,...,n\}\rightarrow\{1,...,n\}$, $U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{\sigma(1)}\cdot...\cdot U_{\sigma(n)}$.
-To do this we will, in fact, prove that
-(*$(*)\ \ \ U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times U_{2}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
-from which the commutativity follows straightforwardly.
-Our main result stands in sharp contrast to the Rudin-Keisler ordering which is known not to be commutative with respect to Fubini product [^2] For example if $U,W$ are non-isomorphic Ramsey ultrafilters then $U\cdot W\not\equiv_{RK}W\cdot U$.
+There is a commutative class of ultrafilter $\mathcal{C}$ which is closed under Fubini sums and includes all Tukey-top ultrafilters, $p$-points, stable ordered union ultrafilters, generic ultrafilters for $P(\omega)/I$ for a wide range of $I$’s.
+[^2] II should be simple, and $I^{\omega}\equiv_{V}I$, so for example $I$ can be $\mathrm{fin}^{\otimes\alpha}$.
+The commutative class described in the previous theorem includes most known ultrafilters on $\omega$, and in fact it is open whether this class consistently excludes an ultrafilter on $\omega$.
+The commutativity of cofinal types stands in sharp contrast to the Rudin-Keisler ordering which is known to be highly non commutative with respect to Fubini product [^3] For example if $U,W$ are non-isomorphic Ramsey ultrafilters then $U\cdot W\not\equiv_{RK}W\cdot U$.
 Just otherwise, by a theorem of Rudin (see for example [20, Thm.
@@
 5.7]) if $U,W$ are $\kappa$-complete ultrafilters on $\kappa$ the $U\cdot W\equiv_{T}W\cdot U$ if and only if $W\equiv_{RK}U^{n}$ for some $n$ or vise versa.
-Recently, Goldberg [18] examined situations of commutativity with respect to several product operations of countable complete ultrafilters.
-As a corollary of our main theorem, we conclude the following Tukey-structural results:
-Theorem.
-The class of ultrafilters $U$ such that $U\cdot U\equiv_{T}U$ is upward closed with respect to the Tukey order.
-The main idea that is used in the proof of the main result, as suggested by the equivalence (*)(*)( * ), is to analyze the cofinal types of ideals and filters connected to a given ultrafilter $U$.
+Recently, Goldberg [18] examined situations of commutativity with respect to several product operations on countably complete ultrafilters.
+The main idea that is used in the proof, is to analyze the cofinal types of ideals and filters connected to a given ultrafilter $U$.
 More specifically, we will exploit the idea of the pseudo-intersection property with respect to $I$ (Definition 2.2) which was introduced in [3] and was used to prove that Milliken-Taylor ultrafilters and generic ultrafilters [^4] For the definition of $I^{\otimes\alpha}$, see the paragraph before Fact 2.11.
 for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfy $U\cdot U\equiv_{T}U$.
-In  §2, we provide a comprehensive study of this property.
-The main result of this section is
+In fact we use the theory developed here to prove some generalization of theorem regarding generic ultrafilters over $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$:
 Theorem.
-Suppose that $\mathcal{A}$ is a discrete set of ultrafilters. Then for each $U\in\mathcal{A}$, $U$ has the $(\cap\mathcal{A})^{*}$-p.i.p.
+Let $I$ be a deterministic $\sigma$-ideal [^5] A $\sigma$-ideal is an ideal such that $P(\omega)/I$ is $\sigma$-complete.
+Then for every $\alpha<\omega_{1}$, and any $V$-generic ultrafilter for $P(\omega^{\alpha})/I^{\otimes\alpha}$, $G\cdot G\equiv_{T}G$.
+In  §2, we provide a comprehensive study of this property. The main result of this section is
+Theorem.
+Suppose that $\mathcal{A}$ is a discrete set of ultrafilters. Then for each $\in\mathcal{A}, UU$ has the $(\bigcap\mathcal{A})^{*}$-p.i.p.
 The property of ideals which will be in frequent use, is the property of being deterministic (Definition 2.22). This property guarantees that whenever $I\subseteq J$, $I\leq_{T}J$.
@@
 Theorem.
-For complete directed ordered set $\mathbb{P}$, $\mathbb{P}$ is uniformly bellow [^5] See definition $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\kappa{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
-By putting more assumptions on the sequence of ultrafilters, we are able to get nicer results. One of them is that under the assumption that
+For complete directed ordered set $\mathbb{P}$, $\mathbb{P}$ is uniformly below  [^7] See Definition 1.11.
+$\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\kappa{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
+By putting more assumptions on the sequence of ultrafilters, we are able to get nicer results. One of them is that under the assumption
 $V_{0}\leq_{T}V_{1}\leq_{T}.$...
 in which case we have that $\sum_{U}V_{n}\equiv_{T}U\times\prod_{n<\omega}V_{n}$.
-From this special case, we can recover the formula $U\cdot V\equiv_{T}U\times V^{\omega}$.
+From this special case, we recover D.
+Milovich’s formula $U\cdot V\equiv_{T}U\times V^{\omega}$ (See Theorem 1.7).
 Another set of assumptions we consider is when for each $n<\omega$, $V_{n}\cdot V_{n}\equiv_{T}V_{n}\geq_{T}V_{n+1}$.
@@
 If $I$ is any ideal such that $I<_{T}\omega^{\omega}$ then $I$ is countably generated.
-this partial order appeared in the work of Molvice [28, Question 4.7], and Dobrinen and Todorcevic who proved the following [14, Thm. 35]:
+On the other hand, among analytic ideals, it is known to be minimal [37](See also [25, Thm 6.6]):
+Theorem 0.2 (Todorcevic).
+Suppose that $I$ is an analytic $p$-ideal, then either $I$ is countable generated or $I\geq_{T}\omega^{\omega}$.
+This was later improved by Solecki and Todorcevic [36, Proposition 4.3] to show that if $I$ is analytic, not locally compact ideal, then $I\geq_{T}\omega^{\omega}$.
+This partial order came up in the work of Milovich who asked [28, Question 4.7] if there is an ultrafilter $U$ over $\omega$ such that $(U,\supseteq)\equiv_{T}\omega^{\omega}$.
+We will observe that this was basically answered in [36, Cor.
+54]:
+Theorem 0.3 (Solecki-Todorcevic).
+Suppose that $D$ is an ordered separable metric space such that the predecessors of each element form a compact set, and $E$ is a basic  [^8] For the definition of basic see [36, §3].
+analytic order such that $D\leq_{T}E$, then $D$ is analytic.
+Later, Dobrinen and Todorcevic contributed a great deal to the understanding of this class, in particular, they proved the following [14, Thm. 35]:
 Theorem 0.4 (Dobrinen-Todorcevic).
@@
     $U\geq_{T}\omega^{\omega}$.
-They then proved that rapid ultrafilters are above $\omega^{\omega}$ and deduced that rapid $p$-points satisfy $U\cdot U\equiv_{T}U$.
-This was lately improved by Benhamou and Dobrinenn [3, Thm 1.18] to the general setup of $I$-p.i.p.
+They also proved that rapid ultrafilters are above $\omega^{\omega}$ and deduced that rapid $p$-points satisfy $U\cdot U\equiv_{T}U$.
+This was lately improved by Benhamou and Dobrinenn [3, Thm 1.18] to the general setup of the $I$-p.i.p.
 Theorem 0.5 (Dobrinen-B.).
@@
 Taking $I=\mathrm{fin}$ reproduces the difficult part from Dobrinen and Todorcevic’s result [^9] Indeed, as pointed out by Dobrinen and Todorcevic, it is easy to see that $U\cdot U\geq_{T}\omega^{\omega}$, so $(1)\Rightarrow(2)$ is straightforward..
-In  §4, we first note that in some sense the assumption that $U$ is $I$-p.i.p in the theorem above is not optimal.
+In  §4, we first note that in some sense the assumption that $U$ has the $I$-p.i.p in the theorem above is not optimal.
 Theorem.
-Let $U$ be any ultrafilter over a set $X$. If $U$ does not have the $I$-p.i.p then $U\geq_{T}\prod_{x\in X}I$.
-In particular, if $U$ is not a $p$-point then automatically $U\geq_{T}\omega^{\omega}$.
-Hence, to understand the class of ultrafilters Tukey above $\omega^{\omega}$, it suffices to restrict our attention to $p$-points.
-Inside the class of $p$-point, it is known to be consistent that there are $p$-point which are not above $\omega^{\omega}$ (see [14]), and the class of rapid p-point is currently the largest subclass of $p$-points known to be above $\omega^{\omega}$.
+Let $U$ be any ultrafilter over a set $X$. If $U$ does not have the $I$-p.i.p then $U\geq_{T}I^{X}$.
+In particular, we get the following corollary:
+Corollary.
+If $U$ is not a $p$-point ultrafilter over $\omega$ then $U\geq_{T}\omega^{\omega}$.
+This last corollary is of the same flavor as Theorems 0.1,0.2.
+Hence if we drop the $p$-point assumption, then we also get $U\geq_{T}\omega^{\omega}$.
+Hence, if we are looking for examples for ideals that are not Tukey above $\omega^{\omega}$ and are not countably generated, we might as well restrict our attention to non-analytic $p$-ideals.
+This is closely related to [14, Question 42].
+Inside the class of $p$-point, it is known to be consistent that there are $p$-points which are not above $\omega^{\omega}$ (see [14]), and the class of rapid p-point is currently the largest subclass of $p$-points known to be above $\omega^{\omega}$.
 In the second part of Section 4, we enlarge this class by introducing the notion of $\alpha$-almost-rapid (Definition 4.12), which is a weakening of rapidness.
@@
   • In §2, we provide a systematic study of the $I$-p.i.p and deterministic ideals.
-  • In §3, we prove our main results.
+  • In §3, we prove our results regarding the commutative classes.
   • In §4 we investigate the class of ultrafilters Tukey above $\omega^{\omega}$.
@@
 Notations
-$[X]^{<\lambda}$ denotes the set of all subsets of $X$ of cardinality less than $\lambda$.
+$X]^{<\lambda}$ denotes the set of all subsets of $X$ of cardinality less than $\lambda$.
 Let $\mathrm{fin}=[\omega]^{<\omega}$, and $\mathrm{FIN}=\mathrm{fin}\setminus\{\emptyset\}$.
@@
 $A\in\sum_{F}G_{x}\text{ if and only if }\{x\in X\mid(A)_{x}\in G_{x}\}\in F$
-[…]
+where $(A)_{x}=\{y\in Y_{x}\mid{\langle}x,y{\rangle}\in A\}$ is the $x^{\text{th}}$-fiber of $A$.
+If for every $x$, $G_{x}=G$ for some fixed $V$ over a set $Y$, then $F\cdot G$ is defined as $\sum_{F}G$, which is a filter over $X\times Y$.
+$F^{2}$ denotes the filter $F\cdot F$ over $X\times X$.
 We distinguish here between $F\cdot G$ and $F\times G$ which is the cartesian product of $F$ and $G$ with the order defined pointwise [^10] There are papers which consider the filter $\{A\times B\mid A\in F,B\in G\}$ and denote it by $F\times G$, this filter will not be considered in this paper so there is no risk of confusion..
@@
 This is the case when $\mathbb{P}_{i}=U_{i}$ is a filter ordered by reversed inclusion of an ideal ordered by inclusion.
-If for every $i\in I$, $\mathbb{P}_{i}=\mathbb{P}$ we simply write $\mathbb{P}^{I}$.
+If for every $\in I, i=\mathbb{P}_{i}$$\mathbb{P}$ we simply write $\mathbb{P}^{I}$.
 1.2. The cofinal type of sums and products
@@
 if said to be coherent if
-$({\dagger})\ \ \text{whenever }Y\subseteq X,\text{ and }C\in U\times\prod_{\alpha\in X}V_{\alpha},\ f_{Y}(\pi_{X,Y}(C))\geq_{\mathbb{P}}f_{X}(C).$
+[…]
 A poset $\mathbb{P}$ is said to be uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if there is a coherent sequence of monotone cofinal maps ${\langle}f_{X}:U\times\prod_{\alpha\in X}V_{\alpha}\rightarrow\mathbb{P}\mid X\in U{\rangle}$.
@@
 Suppose that $\mathbb{P}$ is a complete order.
-Then $\mathbb{P}$ is uniformly bellow $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
+Then $\mathbb{P}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
 Proof.
@@
 $\pi_{\pi_{1}^{\prime\prime}B,\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}X_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}{\rangle})={\langle}\pi_{1}^{\prime\prime}A,{\langle}(A)_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}A{\rangle}{\rangle}$
-[…]
+and that $X_{\alpha}\subseteq(B)_{\alpha}$. It follows by monotonicity of the functions, and by $({\dagger})$ that
 $F(A)=f_{\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}(A)_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}A{\rangle}{\rangle})=f_{\pi_{1}^{\prime\prime}A}(\pi_{\pi_{1}^{\prime\prime}B,\pi_{1}^{\prime\prime}A}({\langle}\pi_{1}^{\prime\prime}A,{\langle}X_{\alpha}\mid\alpha\in\pi_{1}^{\prime\prime}B{\rangle}{\rangle})$
@@
 Lemma 1.13.
-Suppose that $\mathbb{P}$ is complete and for each $X\in U$, $\mathcal{X}_{X}\subseteq U\times\prod_{\alpha\in X}V_{\alpha}$ is such that:
+[…]
   1. (1)
@@
 $f^{*}_{X}(A)=\sup\{f_{X}(B)\mid A\subseteq B\in\mathcal{X}_{X}\}.$
-[…]
+Note that if $B^{\prime}\in\mathcal{X}_{X}$ is such that $B^{\prime}\subseteq A$, then the set $\{f_{X}(B)\mid A\subseteq B\in\mathcal{X}_{X}\}$ is bound in $\mathbb{P}$ by $f_{X}(B^{\prime})$ (as $f_{X}$ is monotone).
+Hence $f^{*}_{X}(A)$ is well defined by completeness of $\mathbb{P}$.
+It is straightforward that Since $f_{X}$ is monotone cofinal, $f^{*}_{X}$ is monotone cofinal.
+To see $({\dagger})$, suppose that $Y\subseteq X$, and $C\in U\times\prod_{\alpha<\lambda}V_{\alpha}$, then for every $C\subseteq B\in\mathcal{X}_{X}$, then by $(3)$ $\pi_{X,Y}(C)\subseteq\pi_{X,Y}(B)\in\mathcal{X}_{X}$ and $f_{Y}(\pi_{X,Y}(B))\geq_{\mathbb{P}}f_{X}(B)$.
+If follows that
 $f^{*}_{X}(C)=\sup\{f_{X}(B)\mid C\subseteq B\in\mathcal{X}_{X}\}\leq\sup\{f_{Y}(\pi_{X,Y}(B))\mid C\leq_{\mathbb{P}}B\in\mathcal{X}_{X}\}$
@@
 $f_{X}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle})={\langle}f_{X(\alpha)}(A_{X(\alpha)})\mid\alpha<\lambda{\rangle}\leq_{\mathbb{P}^{\omega}}{\langle}f_{Y(\alpha)}(A_{Y(\alpha)})\mid\alpha<\lambda{\rangle}$
-$=f_{Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in Y{\rangle}{\rangle})=f_{Y}(\pi_{X,Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}))$
+$f_{Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in Y{\rangle}{\rangle})=f_{Y}(\pi_{X,Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}))$
 Hence by Lemma 1.13 $\mathbb{P}^{\lambda}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ and by Theorem 1.12, $\mathbb{P}^{\lambda}\leq_{T}\sum_{U}V_{\alpha}$. ∎
@@
 $f_{\sigma_{X,X_{0}}^{-1}(n),n}(B_{\sigma_{X,X_{0}}^{-1}(n)})\subseteq B_{n},$
-[…]
+and therefore $f_{X}$ is cofinal.
+Similarly, to see $({\dagger})$, if $Y\subseteq X\subseteq X_{0}$, and ${\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}\in\mathcal{X}_{X}$, then $\sigma_{Y,X_{0}}^{-1}(n)\geq\sigma_{X,X_{0}}^{-1}(n)$, and therefore, for every $n<\omega$,
 $f_{\sigma_{Y,X_{0}}^{-1}(n),n}((A)_{\sigma_{Y,X_{0}}^{-1}(n)})=f_{\sigma_{X,X_{0}}^{-1}(n),n}(f_{\sigma_{Y,X_{0}}^{-1}(n),\sigma_{X,X_{0}}^{-1}(n)}(A_{\sigma_{Y,X_{0}}^{-1}(n)}))\subseteq$
@@
 Let $G\subseteq\mathbb{P}$ be generic over $V$.
-Then in $V[G]$ there is a sequence of ultrafilters $V_{n}$, such that $V_{0}>_{T}V_{1}>_{T}V_{2}...$ and $V_{n}\cdot V_{n}\equiv V_{n}$.
+Then in $V[G]$ there is a sequence of ultrafilters $V_{n}$, such that $V_{0}>_{T}V_{1}>_{T}V_{2}$...
+and $V_{n}\cdot V_{n}\equiv V_{n}$.
 Proof.
@@
 2.1. The pseudo intersection property relative to a set.
-Given set $\mathcal{F}\subseteq P(X)$, we denote by *=$\mathcal{F}^{*}=\{X\setminus A\mid A\in\mathcal{F}\}$.
+Given set $\mathcal{F}\subseteq P(X)$, we denote by $\mathcal{F}^{*}=\{X\setminus A\mid A\in\mathcal{F}\}$.
 When $\mathcal{F}$ is a filter, $\mathcal{F}^{*}$ is an ideal which we call the dual ideal, and when $\mathcal{I}$ is an ideal $\mathcal{I}^{*}$ is a filter which we call the dual filter.
@@
   2. (2)
-[…]
+    $F$ has $\{\emptyset\}$-p.i.p if and only if $F$ is $\sigma$-complete (and therefore, if $F$ is on $\omega$, then it is principal).
 Benhamou and Dobrinen proved the following:
@@
 Let $F$ be any filter. Denoted by $U_{F}\subseteq P(F^{*})$ the set generated by all $T$’s such that $F$ has the $T$-p.i.p, namely,
-[…]
+$U_{F}=\{S\in P(F^{*})\mid\exists T\subseteq S\text{ downwards closed }F\text{ has }T\text{-p.i.p}\}.$
 Then $U_{F}$ is a filter over $F^{*}$.
@@
 If $X\in\bigcap_{n<\omega}U_{n}$, we let $X\cap A_{n}=B_{n}\in U_{n}$ Then $X=X\cap\omega=X\cap(\bigcup_{n<\omega}A_{n})=\bigcup_{n<\omega}X\cap A_{n}=f({\langle}B_{n}\mid n<\omega{\rangle})$.
-To see it is unbounded, suppose that $f^{\prime\prime}\mathcal{A}$ is bounded by $B\in\bigcap_{n<\omega}U_{n}$, then for every $A\in\mathcal{A}$, $A\cap A_{n}\supseteq B\cap A_{n}$.
+To see it is unbounded, suppose that $f^{\prime\prime}\mathcal{A}$ is bounded by $B\in\bigcap_{n<\omega}U_{n}$, then for every $\in\mathcal{A}, AA$$\cap A_{n}\supseteq B\cap A_{n}$.
 and therefore ${\langle}B\cap A_{n}\mid<\omega{\rangle}$ would bound $\mathcal{A}$.
@@
 $\pi^{-1}[\bigcup_{a\in\mathcal{A}}Y\setminus(\pi^{\prime\prime}[X\setminus a])]=\pi^{-1}[Y\setminus(\bigcap_{a\in\mathcal{A}}(\pi^{\prime\prime}[X\setminus a]))]=X\setminus\bigcap_{a\in\mathcal{A}}\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]$
-$=\bigcup_{a\in\mathcal{A}}X\setminus\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]=\bigcup_{a\in\mathcal{A}}a$
+$\bigcup_{a\in\mathcal{A}}X\setminus\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]=\bigcup_{a\in\mathcal{A}}a$
 The last inclusion holds as for each $a$, $X\setminus a=\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]$ as $\pi$ is one-to-one.
@@
 Corollary 2.27.
-Suppose that $I^{\omega}\equiv_{T}I$, $I$ is deterministic. Then for every ultrafilter $U$ such that $I\subseteq U^{*}$ and $U$ satisfies the $I$-p.i.p, $U\cdot U\equiv_{T}U$.
+Suppose that $^{\omega}\equiv_{T}I, II$ is deterministic. Then for every ultrafilter $U$ such that $I\subseteq U^{*}$ and $U$ satisfies the $I$-p.i.p, $U\cdot U\equiv_{T}U$.
 Proof.
@@
 Proposition 2.29.
-Suppose that $\mathrm{fin}\subseteq I$ be any ideal over $\omega$, and ${\langle}J_{n}\mid n<\omega{\rangle}$ is a sequence of deterministic ideals over $\omega$ such that for every $n<\omega$, $J_{n+1}\geq_{T}J_{n}$.
+Suppose that $\mathrm{fin}\subseteq I$ be a deterministic ideal over $\omega$, and ${\langle}J_{n}\mid n<\omega{\rangle}$ is a sequence of deterministic ideals over $\omega$ such that for every $n<\omega$, $J_{n+1}\geq_{T}J_{n}$.
 Then $\sum_{I}J_{n}$ is deterministic.
 Proof.
-Let $\mathcal{B}_{n}\subseteq J_{n}$ witness that $J_{n}$ is deterministic.
+Let $\mathcal{B}_{n}\subseteq J_{n}$ witness that $J_{n}$ is deterministic, and $\mathcal{D}$ witness that $I$ is deterministic.
 Let ${\langle}f_{m,n}:J_{n}\rightarrow J_{m}\mid n\leq m<\omega{\rangle}$ be a sequence of unbounded maps.
 Denote by $\mathcal{B}$ the set of all sequences $\vec{b}={\langle}b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n<m<\omega$, $b_{m}\supseteq f_{m,n}(b_{n})$.
-For $\vec{b}\in\mathcal{B}$ and $A\in I$ defined
+For $\vec{b}\in\mathcal{B}$ and $A\in\mathcal{D}$ defined
 $C_{A,\vec{b}}=\bigcup_{n\in A}\{n\}\times\omega\cup\bigcup_{n\notin A}\{n\}\times b_{n}.$
-Let us show that $\mathcal{C}=\{C_{A,\vec{b}}\mid A\in I,\vec{b}\in\mathcal{B}\}$ is cofinal in $\sum_{I}J_{n}$.
+Let us show that $\mathcal{C}=\{C_{A,\vec{b}}\mid A\in\mathcal{D},\vec{b}\in\mathcal{B}\}$ is cofinal in $\sum_{I}J_{n}$.
 Let $Z\in\sum_{I}J_{n}$, and $A=\{n\mid(Z)_{n}\notin J_{n}\}$.
-By definition of $\sum_{I}J_{n}$, $A\in I$.
+By definition of $\sum_{I}J_{n}$, $A\in I$, so there is $D\in\mathcal{D}$ such that $A\subseteq D$.
 Construct an increasing sequence $\langle b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n\notin A$, $(Z)_{n}\subseteq b_{n}$ and for every $n<m<\omega$, $f_{m,n}(b_{n})\subseteq b_{m}$.
@@
 These are finitely many sets in $J_{n}$ and therefore, since $\mathcal{B}_{n}$ is a cofinal set in $J_{n}$, we can find a single $b_{n}\in\mathcal{B}_{n}$ including the union of these sets.
-It follows that for every $n$, $(Z)_{n}\subseteq(C_{A,{\vec{b}}})_{n}$ and therefore $Z\subseteq C_{A,\vec{b}}$.
+It follows that for every $n$, $(Z)_{n}\subseteq(C_{D,{\vec{b}}})_{n}$ and therefore $Z\subseteq C_{D,\vec{b}}$.
 Let us prove that $\mathcal{C}$ witnesses that $\sum_{I}J_{n}$ is deterministic.
@@
 Note that $(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}$.
-If there is $i_{0}\in T$ such that $n\in X_{i_{0}}$, then $(C_{X_{i_{0}},\vec{b}^{i_{0}}})_{n}=\omega$ and in particular $\omega=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}$.
-Otherwise, consider $n\in A$ such that for every $i\in T$, $n\notin X_{i}$.
-Then $(C_{X_{i},\vec{b}^{i}})_{n}\in\mathcal{B}_{n}$ for every $i\in T$ and $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\notin J_{n}$.
-Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J^{*}_{n}$.
-Fix any $n_{0}\in A$, which exists  [^17] If $I$ is principle then we pick $n_{0}$ such that $\{n_{0}\}\in I^{+}$.
-as $A\notin I$.
+Let us split into cases: If $A\subseteq\bigcup_{i\in T}X_{i}$, then $\bigcup_{i\in T}X_{i}\notin I$ and since $\{X_{i}\mid i\in T\}\subseteq\mathcal{D}$, $\bigcup_{i\in T}X_{i}\in I^{*}$.
+Now for every $n\in\bigcup_{i\in T}X_{i}$, there is $i_{0}\in T$ such that $(C_{X_{i_{0}},\vec{b}^{i_{0}}})_{n}=\omega$ and in particular $\omega=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}$.
+We conclude that $\{n<\omega\mid(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}\}\supseteq\bigcup_{i\in T}X_{i}\in I^{*}$.
+Hence $\bigcup_{i\in T}C_{i}\in(\sum_{I}J_{n})^{*}$.
+Otherwise, consider $n_{0}\in A$ such that for every $i\in T$, $n_{0}\notin X_{i}$.
+Then $(C_{X_{i},\vec{b}^{i}})_{n_{0}}\in\mathcal{B}_{n_{0}}$ for every $i\in T$ and $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n_{0}}\notin J_{n_{0}}$.
+Since $\mathcal{B}_{n_{0}}$ witnesses that $J_{n_{0}}$ is deterministic, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n_{0}}\in J^{*}_{n_{0}}$.
 For every $n\geq n_{0}$, either there is $i\in T$ such that $n\in X_{i}$, and as we have seen, $\cup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}=\omega\in J^{*}_{n}$.
 Otherwise, for every $i\in T$, $(C_{X_{i},\vec{b}^{i}})_{n}=b^{i}_{n}\in\mathcal{B}_{n}$, and by the assumption, $f_{n,n_{0}}(b^{i}_{n_{0}})\subseteq b^{i}_{n}$.
-Since $n_{0}\in A$, $\bigcup_{i\in T}b^{i}_{n_{0}}\notin J_{n_{0}}$, and since $f_{n,n_{0}}$ is unbounded, $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\notin J_{n}$.
+Since $\bigcup_{i\in T}b^{i}_{n_{0}}\notin J_{n_{0}}$, and $f_{n,n_{0}}$ is unbounded, $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\notin J_{n}$.
 Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, it follows that $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\in J^{*}_{n}$.
@@
 Corollary 2.30.
-Suppose that $I,J$ are ideals over $\omega$ such that $J$ is deterministic. Then $I\otimes J$ is deterministic.
+Suppose that $I,J$ are deterministic ideals over $\omega$. Then $I\otimes J$ is deterministic.
+Recall that for an ideal $I$ on $\omega$ we define recursively the ideal $I^{\otimes\alpha}$ on $\omega^{\alpha}$ by setting $I^{1}=I$, $I^{\otimes(\alpha+1)}=I\otimes I^{\otimes\alpha}$ on $\omega^{\alpha+1}=\omega\times\omega^{\alpha}$.
+For limit $\alpha<\omega_{1}$, fix a cofinal sequence ${\langle}\alpha_{n}\mid n<\omega{\rangle}$ in $\alpha$, then $I^{\otimes\alpha}=\sum_{I}I^{\otimes\alpha_{n}}$ is an ideal on $\bigcup_{n<\omega}\{n\}\times\omega^{\alpha_{n}}$.
+It is clear that $\alpha\leq\beta$ then $I_{\alpha}\leq_{RK}I_{\beta}$ (see for example [3, Lemma 3.2]).
 Corollary 2.31.
-For every $I$, $I\otimes\mathrm{fin}$ is deterministic.
+If $I$ is deterministic then for every $\alpha<\omega_{1}$, $I^{\otimes\alpha}$ is deterministic. In particular $\mathrm{fin}^{\otimes\alpha}$ is deterministic for every every $\alpha<\omega_{1}$.
+Proof.
+By induction on $\alpha$.
+For $\alpha=1$ II is deterministic by assumption.
+Suppose that this was true for $\alpha$, then $I^{\alpha+1}=I\otimes I^{\otimes\alpha}$ which is deterministic from the previous corollary.
+Similarily, at limit steps we just use the induction hypothesis and Proposition 2.29.
+In [3] it was noticed that a generic ultrafilter for $P(X)/I$ where $I$ is a $\sigma$-ideal satisfies the $I$-p.i.p.
+now together with Corollary 2.27, we recover the result from [3] in our abstract settings.
+Fact 2.32.
+For every $\alpha\geq 2$, $I^{\otimes\alpha}\equiv_{T}I\otimes I\equiv_{T}I^{\omega}$.
+Proof.
+By induction on $\alpha\geq 2$. For $\alpha=2$,
+$(I\otimes I)^{\omega}\equiv_{T}(I^{\omega})^{\omega}\simeq I^{\omega}\equiv_{T}I\otimes I.$
+For successor $\alpha+1$, we use that fact that $I\otimes I\leq_{T}I^{\otimes\alpha}$ to see that
+$I^{\otimes(\alpha+1)}\equiv_{T}I\otimes I^{\otimes\alpha}\equiv_{T}I\otimes(I\otimes I)\equiv_{T}I\otimes I$
+At limit steps we have that
+$I^{\otimes\alpha}\equiv_{T}\sum_{I}I^{\otimes\alpha}\equiv_{T}\sum_{I}I\times I\equiv_{T}I\otimes(I\otimes I)\equiv_{T}I\otimes I$
+In particular, $(I^{\otimes\alpha})^{\omega}\equiv_{T}I^{\otimes\alpha}$.
 Corollary 2.33.
-For any non-principal ultrafilter $U$ such that $U\equiv_{T}U\cdot U$ and for every $I\subseteq U^{*}$, $I^{\omega}\leq_{T}U$.
+Let $I$ be a deterministic $\sigma$-ideal. Then for every $2\leq\alpha<\omega_{1}$, a generic ultrafilter $G$ for $P(\omega^{\alpha})/I^{\otimes\alpha}$ satisfies $G\cdot G\equiv_{T}G$.
 Proof.
-$I\leq_{T}I\otimes\mathrm{fin}\leq_{T}U\cdot U$. Hence by Corollary 1.8
-$I^{\omega}\leq_{T}(U\cdot U)^{\omega}\equiv_{T}(U\cdot U)\cdot(U\cdot U)\equiv_{T}U\cdot U\equiv_{T}U.$
-3. Proof of the main result and some corollaries
-Let us turn to the proof of the main result which appears in Theorem 3.2 below.
-Corollary 3.1.
-For any non-principal ultrafilter $U,V$ over $\omega$, $V\cdot U,\ U\cdot V\geq_{T}(U^{*}\cap V^{*})^{\omega}$
+$G$ has the $I^{\otimes\alpha}$-p.i.p, and by Corollary 2.31 $I^{\otimes\alpha}$ is deterministic.
+Also by the previous fact, $(I^{\otimes\alpha})^{\omega}\equiv_{T}I^{\otimes\alpha}$ .
+Thus by Corollary 2.27 $G\cdot G\equiv_{T}G$.
+∎
+For $\alpha=1$ we can also say something:
+Proposition 2.34.
+Let $I$ be a simple $\sigma$-ideal. If $G\subseteq P(\omega)/I$ is $V$-generic then $G\cdot G\equiv_{T}G\times I^{\omega}$. So in particular if $I^{\omega}\equiv_{T}I$, $G\cdot G\equiv_{T}G$
 Proof.
-We have that $U^{*}\cap V^{*}\subseteq V^{*}$ and $\mathrm{fin}\subseteq V^{*}$.
-We conclude that $(U^{*}\cap V^{*})\otimes\mathrm{fin}\subseteq(V\cdot V)^{*}$.
-By Corollary 2.31, $(U^{*}\cap V^{*})\otimes\mathrm{fin}$ is deterministic, hence
-$U^{*}\cap V^{*}\leq_{T}(U^{*}\cap V^{*})\otimes\mathrm{fin}\leq_{T}V\cdot V$
-By Milovich’s Corollary 1.8 and Theorem 1.7
-$U\cdot V\equiv_{T}U\cdot(V\cdot V)\equiv_{T}U\times(V\cdot V)^{\omega}\geq_{T}(U^{*}\cap V^{*})^{\omega}.$
-The proof that $V\cdot U\geq_{T}(U^{*}\cap V^{*})^{\omega}$ is symmetric. ∎
-Theorem 3.2.
-Suppose that $U,V$ are ultrafilters on $\omega$, then
-$U\cdot V\equiv_{T}U\times V\times(U\cap V)^{\omega}.$
-Hence $U\cdot V\equiv_{T}V\cdot U$.
-Proof.
-By the previous corollary, $(U\cap V)^{\omega}\equiv_{T}(U^{*}\cap V^{*})^{\omega}\leq_{T}U\cdot V$.
-Since $U\times V\times(U\cap V)^{\omega}$ is the least upper bound in the Tukey order of $U,V,(U\cap V)^{\omega}$, we have
-$U\times V\times(U\cap V)^{\omega}\leq_{T}U\cdot V$
-For the other direction, recall that by Theorem 2.13 $V$ has the $(U\cap V)^{*}$-p.i.p and therefore by Proposition 2.5
-$U\cdot V\equiv_{T}U\times V^{\omega}\leq_{T}U\times V\times(U^{*}\cap V^{*})^{\omega}\equiv_{T}U\times V\times(U\cap V)^{\omega}.$
+Since $G$ has the $I$-p.i.p, $G\cdot G\leq_{T}G\times I^{\omega}$.
+In the other direction, we known that $I^{*}\subseteq G$ and since $I$ is simple, $I\leq_{T}G$.
+Hence $I^{\omega}\leq_{T}G^{\omega}\equiv_{T}G\cdot G$.
+We conclude that $G\times I^{\omega}\leq G\cdot G$.
+∎
+3. Commutativity of cofinal types
+Definition 3.1.
+We say that a class $\mathcal{C}$ of ultrafilters on $\omega$ is a commutative class if for every $U,W\in\mathcal{C}$, $U\cdot W\equiv_{T}W\cdot U$.
+Example 3.2.
+By Dobrinen and Todorcevic, the class $\mathcal{C}_{RP}$ of rapid $p$-points is a commutative class and if $W$ is a rapid $p$-point then $U\cdot W\equiv_{T}U\times W$.
+Example 3.3.
+By Milovich [29], the class $\mathcal{C}_{P}$ of $p$-points is a commutative class since if $W$ is a $p$-point, then $U\cdot W\equiv_{T}U\times W\times\omega^{\omega}$.
+Fact 3.4.
+For any three ultrafilters $U,W,Z$, $U\cdot(W\cdot Z)\equiv_{RK}(U\cdot W)\cdot Z$
+From the above associativity, we get the following:
+Proposition 3.5.
+If $\mathcal{C}$ is a commutative class then so is $\overline{\mathcal{C}}=\{U_{1}\cdot U_{2}...\cdot U_{n}\mid U_{1},...,U_{n}\in\mathcal{C}\}$.
+One use of commutative classes relates to the character of an ultrafilter.
 Recall that for an ultrafilter $U$, $Ch(U)$ denoted the minimal size of a cofinal subset of $U$.
 It is clear that if $U\equiv_{T}U^{\prime}$ then $Ch(U)=Ch(U^{\prime})$.
-As a corollary of theorem 3.2, we get:
+Hence we conclude the following:
 Corollary 3.6.
-For any two ultrafilters, $Ch(U\cdot V)=Ch(V\cdot U)$.
-It is possible to prove by induction now that the product of $n$-ultrafilters commute, but we would like to get the exact cofinal type of such product. We will need the following facts:
-Fact 3.4.
-  1. (1)
-    $U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times(U_{2})^{\omega}...\times(U_{n})^{\omega}$.
-  2. (2)
-    $(\mathbb{P}^{\omega})^{\omega}\simeq\mathbb{P}^{\omega}$.
+Let $\mathcal{C}$ be a commutative class, then for any two ultrafilters $U,V\in\mathcal{C}$, $Ch(U\cdot V)=Ch(V\cdot U)$.
+Definition 3.7.
+Let $D$ be a cofinal type, define $\mathcal{E}_{D}=\{U\mid U\cdot U\equiv_{T}U\times D\}$.
+For example, $\mathcal{E}_{\{0\}}=\mathcal{E}_{\omega}=\{U\mid U\cdot U\equiv_{T}U\}$. Also $\mathcal{C}_{P}\subseteq\mathcal{E}_{\omega^{\omega}}$.
+Claim 3.8.
+The class $\mathcal{E}_{\{0\}}\subseteq\mathcal{E}_{\omega^{\omega}}$.
 Proof.
-$(1)$ is just by induction using Milovich Theorem 1.7. For $(2)$, just decompose $\omega$ into infinitely many infinite sets. See for example [3, Fact 2.4]. ∎
-Corollary 3.5.
-For every ultrafilters $U_{1},...,U_{n}$,
-$U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times U_{2}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
+Just note that $\omega^{\omega}\leq_{T}U\cdot U$ for every uniform ultrafilter $U$ and therefore if $U\cdot U\equiv_{T}U$ then $U\equiv_{T}U\times\omega^{\omega}$ and in particular $U\cdot U\equiv_{T}U\times\omega^{\omega}$.
+∎
+Proposition 3.9.
+For every cofinal type $D$, $\mathcal{E}_{D}$ is a commutative class
 Proof.
-Following the proof of Theorem 3.2, $U_{1}\cap...\cap U_{n}\subseteq U_{n}$ and we can prove similarly that $(U_{1}\cap....\cap U_{n})^{\omega}\leq_{T}U_{n}\cdot U_{n}\leq_{T}U_{1}\cdot...\cdot U_{n}$.
-We conclude that
-$U_{1}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}\leq_{T}U_{1}\cdot...\cdot U_{n}$
-In the other direction, each $U_{i}$ has the $(U_{1}\cap...\cap U_{n})^{*}$-p.i.p and therefore
-$U_{1}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times(U_{2})^{\omega}\times(U_{3})^{\omega}\times...\times(U_{n})^{\omega}\leq_{T}$
-$\leq_{T}U_{1}\times U_{2}\times(U_{1}\cap....\cap U_{n})^{\omega}\times U_{3}\times(U_{1}\cap...\cap U_{n})^{\omega}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
-$\leq_{T}U_{1}\times...\times U_{n}\times((U_{1}\cap...\cap U_{n})^{\omega})^{\omega}\equiv_{T}U_{1}\times....\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
-Corollary 3.6.
-Suppose that $U\cdot U\equiv_{T}U$ then for every $V\geq_{T}U$, $V\cdot V\equiv_{T}V$.
-Namely the class of ultrafilters $U$ such that $U\cdot U\equiv_{T}U$ is upward closed and the class of ultrafilters $V$ such that $V\cdot V>_{T}V$ is downward closed.
+Indeed, for any $U,V\in\mathcal{E}_{D}$, by Theorem 1.7
+$V\cdot U\equiv_{T}V\times U^{\omega}\equiv_{T}V\times U\cdot U\equiv_{T}V\times U\times D$
+Since the formula above is symmetric for $V,U$, we conclude that $U\cdot V\equiv_{T}V\cdot U$. ∎
+Proposition 3.10.
+The commutative class $\mathcal{E}_{\omega^{\omega}}$ contains all Tukey-top ultrafilters, $p$-points, stable ordered union ultrafilters and generic ultrafilters for $P(\omega)/I$ where $I^{\omega}\equiv_{T}I$ is a simple $\sigma$-ideal
 Proof.
-Otherwise, $V<_{T}V\cdot V$ and
-$U\cdot V\equiv_{T}U\times V^{\omega}\equiv_{T}V^{\omega}\equiv_{T}V\cdot V>_{T}V\equiv_{T}V\times U\equiv_{T}V\times U^{\omega}\equiv_{T}V\cdot U,$
-contradiction. ∎
-Corollary 3.7.
-For every $\alpha$, $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
+Clearly Tukey top ultrafilters are in any $\mathcal{E}_{D}$, for $D$ of size at most continuum.
+$p$-points are in $\mathcal{E}_{\omega^{\omega}}$ as we have seen by Milovich [29], stable ordered union and generic ultrafilters for $P(\omega^{\alpha})/I$ are in fact in $\mathcal{E}_{\{0\}}$, by 2.27.
+∎
+Theorem 3.11.
+$\mathcal{E}_{\omega^{\omega}}$ is closed under Fubini sums.
 Proof.
-The base case $\mathrm{fin}^{\otimes 1}=\mathrm{fin}$ which is deterministic.
-Suppose this is true for $\alpha$, then $\mathrm{fin}^{\otimes\alpha+1}=\mathrm{fin}\cdot\mathrm{fin}^{\otimes\alpha}$ and by induction hypothesis and Corollary 2.30 $\mathrm{fin}^{\otimes\alpha+1}$ is deterministic.
-At limit step, $\mathrm{fin}^{\otimes\alpha}=\sum_{\mathrm{fin}}\mathrm{fin}^{\otimes\alpha_{n}}$.
-By it is known that $\mathrm{fin}^{\alpha_{n}}$ is increasing in the Tukey order (even in the Rudin-Keisler order, see for example [3, Lemma 3.2]) and by induction hypothesis are all deterministic.
-Therefore, by Proposition 2.29 $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
-In [3] it was noticed that a generic ultrafilter for $P(X)/I$ where $I$ is a $\sigma$-ideal satisfies the $I$-p.i.p.
-now together with Corollary 2.27, we recover the result from [3] in our abstract settings:
-Corollary 3.8.
-For every $\alpha$, a generic ultrafilter $G$ for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfies $G\cdot G\equiv_{T}G$.
+Suppose that $\{W,W_{0},W_{1},...\}\subseteq\mathcal{E}_{\omega^{\omega}}$. We need to prove that
+$\sum_{W}W_{n}\cdot\sum_{W}W_{n}\equiv_{T}\sum_{W}W_{n}\times\omega^{\omega}.$
+Note that $\sum_{W}W_{n}\geq_{T}\omega^{\omega}$, so we will end up getting $\sum_{W}W_{n}\cdot\sum_{W}W_{n}\equiv_{T}\sum_{W}W_{n}$. It is not hard to see that
+$\sum_{W}W_{n}\cdot\sum_{W}W_{n}\geq_{T}\sum_{W}W_{n}\times\omega^{\omega}.$
+For the other direction, recall that $\sum_{W}W_{n}\cdot\sum_{W}W_{n}\equiv_{T}(\sum_{W}W_{n})^{\omega}$.
+Let us prove that $(\sum_{W}W_{n})^{\omega}$ is a uniformly below $\mathcal{B}(W,{\langle}W_{n}\mid n<\omega{\rangle})$.
+Since $(\sum_{W}W_{n})^{\omega}$ is complete, Theorem 1.12 can be applied that get $\sum_{W}W_{n}\geq_{T}(\sum_{W}W_{n})^{\omega}$.
+Let $f:W\times\omega^{\omega}\rightarrow W^{\omega}$ and $f_{n}:W_{n}\times\omega^{\omega}\rightarrow W_{n}^{\omega}$ be monotone and cofinal maps, which exists by the assumption that $W,W_{n}\in\mathcal{E}_{\omega^{\omega}}$.
+We need to define a sequence of monotone and cofinal maps ${\langle}g_{X}:W\times\prod_{n\in X}W_{n}\rightarrow(\sum_{W}W_{n})^{\omega}\mid X\in W{\rangle}$ such that $(\dagger)$ of definition 1.11 holds.
+Let ${\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}\in W\times\prod_{n\in X}W_{n}$, by lemma 1.13, we may restrict ourselves to sequences satisfying that
+[…]
+The first step is to produce $\omega$-many functions in $\omega^{\omega}$ which are going to be the inputs of $f_{n}$’s.
+Fix a partition of $\omega=\biguplus_{l<\omega}Z_{l}$ such that each $Z_{l}$ is infinite.
+Recall that for a set $C$ on natural numbers, $C(r)$ denotes the $r^{\text{th}}$ element of $C$ in its increasing enumeration.
+Define $\varphi_{X,i}$ by induction on $i$.
+For $i=0$, $\varphi_{X,0}(k)=X(Z_{0}(k))$.
+Inductively, $\varphi_{X,{i+1}}(k)=\max(\varphi_{X,i}(k),X(Z_{i+1}(k)))$.
+Claim 3.12.
+If $Y\subseteq X$, then for every $i$, $\varphi_{X,i}\leq\varphi_{Y,i}$.
 Proof.
-$G$ has the $\mathrm{fin}^{\otimes\alpha}$-p.i.p, and by Corollary 3.7 $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
-Also note that $(\mathrm{fin}^{\otimes\alpha})^{\omega}\equiv_{T}\mathrm{fin}^{\otimes\alpha}$ (see [3, Thm.
-3.3 & Fact 2.4]).
-Thus by Corollary 2.27 $G\cdot G\equiv_{T}G$.
+Clearly, for every $m<\omega$, $X(m)\leq Y(m)$.
+So by definition, $\varphi_{X,0}(k)\leq\varphi_{Y,0}$.
+Suppose this was true for $i$, and let $k<\omega$, then by the induction hypothesis and our first observation,
+$\max(\varphi_{X,i}(k),X(Z_{i+1}(k)))\geq\max(\varphi_{Y,i}(k),Y(Z_{i+1}(k)).$
+The $i$’s function we will use is $h^{i}_{{\langle}A_{n}\mid n\in X{\rangle}}(k)=\min(A_{\varphi_{X,i}(k)})$.
+This is well defined as by definition of $\varphi_{X,i}$, $\varphi_{X,i}(k)\in X$.
+If $Y\subseteq X$, then by the claim $\varphi_{X,i}(k)\leq\varphi_{Y,i}(k)$, and by $(\star)$, $h^{i}_{{\langle}A_{n}\mid n\in X{\rangle}}(k)\leq h^{i}_{{\langle}A_{n}\mid n\in Y{\rangle}}(k)$.
+Now define $g_{X}({\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle})$ by
+${\langle}\pi_{X}({\langle}f(B,h^{0}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m},{\langle}f_{n}(A_{n},h^{n+1}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m}\mid n\in X{\rangle})\mid m<\omega{\rangle}$
+The above seemingly complicated definition is nothing but the composition of the following quite natural monotone cofinal maps:
+$W\times\prod_{n\in X}W_{n}\overset{(id,{\langle}h^{i}_{*}\mid i<\omega{\rangle})}{\longrightarrow}W\times\prod_{n\in X}W_{n}\times(\omega^{\omega})^{\omega}\rightarrow(W\times\omega^{\omega})\times\prod_{n\in X}(W_{n}\times\omega^{\omega})\overset{(f,{\langle}f_{n}\mid n\in X{\rangle})}{\longrightarrow}$
+$\overset{(f,{\langle}f_{n}\mid n\in X{\rangle})}{\longrightarrow}W^{\omega}\times\prod_{n\in X}W_{n}^{\omega}\longrightarrow(W\times\prod_{n\in X}W_{n})^{\omega}\overset{\pi_{X}^{\omega}}{\longrightarrow}(\sum_{W}W_{n})^{\omega}$
+So $g_{X}$ is clearly monotone cofinal as the composition of such functions.
+To see $(\dagger)$, let $Y\subseteq X$, we ensured that $h^{i}_{{\langle}A_{n}\mid n\in X{\rangle}}\leq h^{i}_{{\langle}A_{n}\mid n\in Y{\rangle}}$.
+Since $f,f_{n}$ are monotone functions, for any $m<\omega$, and any $n\in Y$,
+$f(B,h^{0}_{{\langle}A_{n}\mid n\in Y{\rangle}})_{m}\subseteq f(B,h^{0}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m}$
+and
+$f_{n}(A_{n},h^{n+1}_{{\langle}A_{n}\mid n\in Y{\rangle}})_{m}\subseteq f_{n}(A_{n},h^{n+1}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m}.$
+By definition of $\pi_{X}$ and $\pi_{Y}$, as unions of the relevant sets (see p.
+9) we conclude that $g_{Y}(\pi_{X,Y}({\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}))\geq g_{X}({\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle})$, as wanted.
 ∎
+Given a cofinal type $D$, by a (yet) finer handling of coherent sequences of functions, we can define an analog of condition $(\dagger)$, that ensure we can amalgamate monotone cofinal maps $g_{X}:U\times\prod_{n\in X}V_{n}\times D\rightarrow\mathbb{P}$, to a monotone cofinal map $g:\sum_{U}V_{n}\times D\rightarrow\mathbb{P}$.
+Thus proving an analog of Theorem 1.12 that $\sum_{U}V_{n}\times D$ is the greatest lower bound of $\{U\times\prod_{n\in X}V_{n}\times D\mid X\in U\}$ among uniform lower bounds.
+Then the proof of the previous Theorem generalizes to give the next theorem.
+Since the proofs are essentially the same, we do not include them.
+Theorem 3.13.
+Suppose that $D^{\omega}\equiv_{T}D$ then $\mathcal{E}_{D}$ is closed under Fubini sums.
+We do not know whether the class $\mathcal{E}_{\omega^{\omega}}$ is provably all uniform ultrafilters on $\omega$.
+We conjecture that consistently it is not but this will require to produce new examples of non-Tukey top ultrafilters, as it is unknown whether there is a basically generated ultrafilter which is not in the (transfinite) closure of the $p$-points under Fibini sums, and the only examples for non-Tukey-top non-basically generated ultrafilters are generic ultrafilters for $P(\omega^{\otimes\alpha})/\mathrm{fin}^{\otimes\alpha}$.
 4. On ultrafilters above $I^{\omega}$
-The cofinal type of $\omega^{\omega}$ came up in several early papers [28, 35] regarding the Tukey order on ultrafilters.
+The cofinal type of $\omega^{\omega}$ came up in several papers [28, 36] regarding the Tukey order on general ultrafilters.
+Milovich asked whether there is an ultrafilter $U$ such that $(U,\supseteq)$ is Tukey equivalent to $\omega^{\omega}$.
+Let us point out that a negative answer is a straightforward corollary [^20] Milovich’s question appeared only 4 years after Solecki and Todorcevic’s result.
+of Theorem 0.3:
+Proposition 4.1.
+There is no non-principal ultrafilter $U$ over $\omega$ such that $(U,\supseteq)\equiv_{T}\omega^{\omega}$.
+Proof.
+By Sierpinski [35], a non-principal ultrafilter over $\omega$ is a non-measurable set as a subset of $2^{\omega}$ and in particular non-analytic.
+An ultrafilter $U$ with the topology inherited from $2^{\omega}$ is a separable metric space and the set of predecessors is compact and $\omega^{\omega}$ is a basic analytic order, hence by Theorem 0.3, $(U,\supseteq)\not\leq_{T}\omega^{\omega}$.
+∎
 It turns out that some general problems boil down to being Tukey above $\omega^{\omega}$.
 Such a problem is addressed in Theorem 0.4, which sets up the equivalence for $p$-point ultrafilter $U$, between $U\cdot U\equiv_{T}U$ and $U\geq_{T}\omega^{\omega}$.
-This was genralized in Theorem 0.3 which ensures that for a general ultrafilter $U$, $U\cdot U\equiv_{T}U$ is equivalent to the existence of some $I\subseteq U^{*}$ such that $U$ has the $I$-p.i.p and $U\geq_{T}I^{\omega}$.
+This was generalized in Theorem 0.5 which ensures that for a general ultrafilter $U$, $U\cdot U\equiv_{T}U$ is equivalent to the existence of some $I\subseteq U^{*}$ such that $U$ has the $I$-p.i.p and $U\geq_{T}I^{\omega}$.
 In the first part of this section, we tighten the connection between the $I$-p.i.p and being above $I^{\omega}$ for a deterministic ideal $I$.
 Then, in the second subsection, we shall restrict our attention to $I=\omega$.
-4.1. The case of a deterministic ideal $I$
-There is a slight difference between the type of equivalence in Theorem 0.2 to the equivalence $U\cdot U\equiv_{T}U$ for $p$-points and the general one in 0.3.
+4.1. The case of a deterministic ideal II
+There is a slight difference between the type of equivalence in Theorem 0.4 to the equivalence $U\cdot U\equiv_{T}U$ for $p$-points and the general one in 0.5.
 Indeed, in the latter, the ideal $I$ can vary.
@@
 ∎
+The previous proposition is somewhat different from [3, Thm. 1.18], as here $I$ is a fix simple ideal.
 Our next objective it to study the class of ultrafilters which are Tukey above $I^{\omega}$.
@@
 As observed by Dobrinen and Todorcevic, rapid ultrafilters form a subclass of those which are Tukey above $\omega^{\omega}$.
-By Miller [26], in the Laver model where the Borel conjecture holds181818 i.e.
+By Miller [26], in the Laver model where the Borel conjecture holds212121i.e.
 the model obtained by adding $\omega_{2}$-many Laver reals to a model of $CH$., there are no rapid ultrafilters.
@@
 By strengthening the above definition, we may require that $exp_{\alpha}(f_{X})\geq f$. However, this strengthening turns out to be an equivalent definition.
-Note that 00-almost-rapid is just rapid, and by $(3)$ of the previous lemma, if $\beta\leq\alpha$ then $\beta$-almost-rapid implies $\alpha$-almost-rapid.
+Note that 0-almost-rapid is just rapid, and by $(3)$ of the previous lemma, if $\beta\leq\alpha$ then $\beta$-almost-rapid implies $\alpha$-almost-rapid.
 We call $U$ almost-rapid if it is $1$-almost-rapid.
@@
 Suppose that ${\langle}P_{n}\mid n<\omega{\rangle}$ is not a counterexample for $U$ being rapid, then there is a set $X\in U$ such that $|X\cap P_{n}|\leq n$ for every $n$ and therefore $X\in I$.
-Hence, as long as we have $U\subseteq I^{+}$, $U$ will not be rapid.
+Hence, as long as we have $\subseteq I^{+}, UU$ will not be rapid.
 Note that
-$A\in I^{+}\text{ iff for every }k,\ \text{there is }n_{k}\text{ such that }|A\cap P_{n_{k}}|>kn_{k}.$
+[…]
 Or equivalently, $n\mapsto|A\cap P_{n}|$ is not asymptotically bounded by a linear function of $n$. The following is the key lemma for our construction:
@@
 $|A_{2}\cap P_{n_{2}}\setminus\{0,...,n_{1}+2n_{2}+2\}|>(f(2)+2)(a_{n_{1}+1}+2)n_{2}-(n_{1}+2n_{2}+3)>$
-$>8n_{2}-(3n_{2}+3)=6n_{2}-3>2n_{2}+1$
+$8n_{2}-(3n_{2}+3)=6n_{2}-3>2n_{2}+1$
 So we can find $a_{n_{1}+2},...,a_{n_{1}+1+2n_{2}+1}$ above $n_{1}+1+2n_{2}+1$ (and therefore also above $a_{n_{1}+1}$).
@@
 $|A_{k+1}\cap P_{n_{k+1}}\setminus\{0,...,\sum_{i=1}^{k+1}in_{i}+1\}|>3(k+1)(f(k+1)+1)n_{k+1}-(\sum_{i=1}^{k+1}in_{i}+1)-1$
-$>3(k+1)(f(k+1)+1)n_{k+1}-((k+1)n_{k+1}+1)-(\sum_{i=1}^{k}in_{i}+1)-1$
+$3(k+1)(f(k+1)+1)n_{k+1}-((k+1)n_{k+1}+1)-(\sum_{i=1}^{k}in_{i}+1)-1$
 $(k+1)(3f(k+1)+3)n_{k+1}-2((k+1)n_{k+1}+1)$
-$>(k+1)(3f(k+1)+1)n_{k+1}>(k+1)n_{k+1}+1$
+$(k+1)(3f(k+1)+1)n_{k+1}>(k+1)n_{k+1}+1$
 Hence we can find $(k+1)n_{k+1}+1$-many elements in $A_{k+1}\cap P_{k+1}$ above $\sum_{i=1}^{k+1}in_{i}+1$.
@@
 At limit steps $\delta$ we define $V_{\delta}=\bigcup_{\beta<\delta}V_{\beta}$.
-It is clear that $(1)-(5)$ still holds at limit steps.
+It is clear that (1)(5)(1)-(5) still holds at limit steps.
 At successors, given $V_{\alpha}$, since we have only performed countably many steps so far, there are sets $B_{n}\in V_{\alpha}$ such that $V_{\alpha}=I^{*}[{\langle}B_{n}\mid n<\omega{\rangle}]$ where $B_{n}$ is $\supseteq$-decreasing.
@@
 If $\vec{A}_{\alpha}\not\subseteq V_{\alpha}[X_{\alpha}]$ ignore it.
-Otherwise, enumerate the set $\{B_{n}\cap X_{\alpha}\mid n<\omega\}\cup\{\vec{A}_{\alpha}(n)\mid n<\omega\}\subseteq V_{\alpha}[X_{\alpha}]$ by ${\langle}B^{\prime}_{n}\mid n<\omega{\rangle}$ and let $C_{n}=\cap_{m\leq n}B^{\prime}_{m}$.
+Otherwise, enumerate the set
+$\{B_{n}\cap X_{\alpha}\mid n<\omega\}\cup\{\vec{A}_{\alpha}(n)\mid n<\omega\}\subseteq V_{\alpha}[X_{\alpha}]$
+by ${\langle}B^{\prime}_{n}\mid n<\omega{\rangle}$ and let $C_{n}=\cap_{m\leq n}B^{\prime}_{m}$.
 We apply the previous lemma to the sequence ${\langle}C_{n}\mid n<\omega{\rangle}$, and $\tau_{\alpha}$ to find $A^{*}\subseteq\omega$ such that:
@@
 However, $A^{*}\subseteq^{*}B_{n}\cap X_{\alpha}$, which implies that $A^{*}\in I$, contradicting property $(1)$ above in the choice of $A^{*}$.
-Hence we can define $V_{\alpha+1}=V_{\alpha}[X_{\alpha},A^{*}]$ and $(1)-(5)$ hold.
-This concludes the recursive definition. The ultrafilter witnessing the theorem is defined by $V^{*}=\cup_{\alpha<\omega_{1}}V_{\alpha}$.
+Hence we can define $V_{\alpha+1}=V_{\alpha}[X_{\alpha},A^{*}]$ and (1)(5)(1)-(5) hold.
+This concludes the recursive definition. The ultrafilter witnessing the theorem is defined by $V^{*}=\bigcup_{\alpha<\omega_{1}}V_{\alpha}$.
 Proposition 4.18.
@@
 Is the equivalence of Proposition 4.2 true for every ideal $I$?
-The following addresses the commutativity of cofinal types between ultrafilters on different cardinals.
+The following addresses the possibility of full commutativity of cofinal types between ultrafilters on $\omega$ and on different cardinals.
 Question 5.8.
+Is it provable that for any two ultrafilters $U,V$, $U\cdot V\equiv_{T}V\cdot U$?
+Question 5.9.
 Is it true that for every two ultrafilters $U,V$ on any cardinals, $U\cdot V\equiv_{T}V\cdot U$?
@@
 So the question above is interesting in case we drop the completeness assumption on the ultrafilters.
+the results in § 4 suggest that the Tukey-type of $I^{\omega}$ as an important role in the analysis of non Tukey-top ultrafilters.
+Note that the same argument which worked for $\omega^{\omega}$ in Proposition 4.1 works for $I^{\omega}$, hence leading to the conclusion that if $I$ is an analytic ideal the no non-principal ultrafilter can be Tukey equivalent to $I^{\omega}$.
+The following question is more open-ended:
+Question 5.10.
+How can we force different values of the Tukey-type of $I^{\omega}$, when $I\not\equiv_{T}\omega^{\omega}$?
 Finally, we present a few questions regarding the new class of $\alpha$-almost-rapid ultrafilters.
-Question 5.9.
+Question 5.11.
 Is it true that for every $\alpha<\beta<\omega_{1}$, the class of $\alpha$-almost-rapid ultrafilters is consistently strictly included in the class of $\beta$-almost-rapid ultrafilters?
 We conjecture a positive answer to this question and that similar methods to the one presented in Theorem 4.16 under CH should work.
-Question 5.10.
+Question 5.12.
 Does $\mathfrak{d}=\mathfrak{c}$ imply that there is a $p$-point which is almost-rapid but not rapid?
 Ultimately we would like to understand if such ultrafilters exist in $ZFC$:
-Question 5.11.
+Question 5.13.
 Is it consistent that there are no almost-rapid ultrafilters?
@@
 I would like to thank Gabriel Goldberg for stimulating discussions and for suggesting an example of ultrafilters which eventually led to Proposition 1.21.
-I would also like to express my gratitude to Natasha Dobrinen for introducing me to the subject, a task which took a great amount of her precious time.
+I would also like to thank Slawomir Solecki and Stevo Todorcevic for answering some questions regarding their work on the Tukey order.
+Finally, a special gratitude goes to Natasha Dobrinen for introducing me to the subject, a task which took a great amount of her precious time, and for her comments on early drafts of this paper.